### v1

# Can Few-shot Work in Long-Context? Recycling the Context to Generate Demonstrations

Abstract
Despite recent advancements in Large Language Models (LLMs), their performance on tasks involving long contexts remains sub-optimal. In-Context Learning (ICL) with few-shot examples may be an appealing solution to enhance LLM performance in this scenario; However, naïvely adding ICL examples with long context introduces challenges, including substantial token overhead added for each few-shot example and context mismatch between the demonstrations and the target query. In this work, we propose to automatically generate few-shot examples for long context QA tasks by recycling contexts. Specifically, given a long input context (1-3k tokens) and a query, we generate additional query-output pairs from the given context as few-shot examples, while introducing the context only once. This ensures that the demonstrations are leveraging the same context as the target query while only adding a small number of tokens to the prompt. We further enhance each demonstration by instructing the model to explicitly identify the relevant paragraphs before the answer, which improves performance while providing fine-grained attribution to the answer source. We apply our method on multiple LLMs and obtain substantial improvements on various QA datasets with long context, especially when the answer lies within the middle of the context. Surprisingly, despite introducing only single-hop ICL examples, LLMs also successfully generalize to multi-hop long-context QA using our approach.
[…]
tcb@breakable
Can Few-shot Work in Long-Context?
Recycling the Context to Generate Demonstrations
Arie Cattan1,2thanks: Work done during an internship at Google.  Alon Jacovi2  Alex Fabrikant3  Jonathan Herzig2 Roee Aharoni2Hannah Rashkin3Dror Marcus2 Avinatan Hassidim2Yossi Matias2Idan Szpektor2Avi Caciularu2 1Bar-Ilan University  2Google Research  3Google DeepMind cattana@google.com
1 Introduction
Figure 1: Performance of Gemini Pro (v1.0) on a sample of the Lost-in-the-middle dataset (Liu et al., 2023). The X-axis is the position of the relevant passage in the context. The baseline (blue line) displays a U-shaped curve, performing well only when the relevant passage is at the beginning or end of the input. The oracle (green line) shows significant performance gain when the relevant passage ID is provided in the prompt, showing that the main bottleneck is the identification of supporting evidence(s). DoubleDipper (our method, orange line) flattens this U-shaped trend.
Long contexts are prevalent in various domains, ranging from legal documents and scientific articles to lengthy reports and novels. These may consist of a single extensive document or multiple passages, typically retrieved through specific retrieval mechanisms (e.g., RAG; Lewis et al., 2020).
Yet, while Large Language Models (LLMs) have demonstrated impressive capabilities in a variety of tasks including answering questions requiring one or multiple reasoning steps, they often struggle to answer simple questions when faced with long contexts. Despite substantial engineering efforts (Chen et al., 2023) to extend the context window of LLMs to extremely long inputs (32k and even 1M tokens), these models continue to struggle with much shorter inputs, comprising only a few thousand tokens.
In order to answer questions from long inputs, models should implicitly identify relevant information segments and then reason over these segments to formulate an answer. It has been shown that LLMs struggle when the relevant information is buried in the middle of the context (Liu et al., 2023) or obscured by numerous irrelevant details (Levy et al., 2024). Our analysis (illustrated in Figure 1) identifies the identification of relevant information as a primary performance bottleneck of current models in long contexts.
In this work, we introduce a novel method to enhance the QA performance of LLMs in long input setups (to allow direct comparisons across a wide swath of models, we limit "long context" here to 1-3k tokens). Our approach, termed DoubleDipper, leverages LLMs’ In-Context Learning (ICL) capability and is based on two principles. First, instead of typical ICL, where each few-shot example is standalone with a separate length context and a question-answer (QA) pair, we propose to recycle the given input context and automatically generate few-shot examples from this context. Specifically, we randomly select a few paragraphs from the given input context and generate QA pairs for each passage. These generated QAs serve as demonstration examples and are placed between the input context and the target input question. Figure 2 illustrates the differences between the traditional ICL with few-shot examples and DoubleDipper. Second, we further enrich each ICL example to instruct the model to explicitly identify the paragraph(s) containing the relevant information before generating the answer. This can be regarded as a structured Chain of Thought that incentivizes the model to pinpoint relevant information before reasoning, an essential capability for long-context processing.
Figure 2: Comparison of traditional In-Context-Learning (ICL) and our new method. In traditional ICL (left), each example comprises a possibly lengthy context, accompanied by a query and an answer, typically derived from the training dataset. Conversely, our approach (right) simplifies each example to just a question and an answer, both of which are generated directly from the provided input context.
By generating few-shot demonstrations from various sections of the input context while instructing the model to identify relevant passages, DoubleDipper encourages the model to develop deeper reading comprehension skills specific to the given input evidence. This, in turn, allows the model to answer subsequent queries with higher accuracy. DoubleDipper presents several advantages. In terms of efficiency, since each example does not include its own input context, our method adds to the original prompt a minimal number of tokens, resulting in a substantially cheaper inference than traditional ICL. Additionally, recycling the same context for ICL demonstrations ensures that the few-shot examples refer to the same domain as the input question, thus obviating the need for external retrieval processes. Finally, DoubleDipper generates answers with attribution to relevant paragraphs, improving the model’s lookup ability and offering transparency, which substantially simplifies human evaluation (Slobodkin et al., 2024).
We applied DoubleDipper to a variety of LLMs, both commercial (Gemini Pro and Gemini Ultra; Team et al., 2023) and open-source (Llama (Touvron et al., 2023), Mistral (Jiang et al., 2023) and Gemma (Team et al., 2024)), and evaluate it on various QA datasets with long inputs, including common multi-hop QA datasets. Our experiments demonstrate that with only 3 self-generated few-shot examples, DoubleDipper consistently outperforms the baseline on our evaluation set. In addition, for some models, DoubleDipper enhances the robustness to the position of the relevant information within the text. Interestingly, while our few-shot examples focus on single-paragraph answers, DoubleDipper generalizes well to multi-hop QAs where the answer requires information from multiple passages.
2 Background
Challenges in Long Context for Language Modeling.
LLMs have been well-documented to struggle when input length grows (An et al., 2023), and especially so when it exceeds input lengths seen during training (Anil et al., 2022). Various methods have been proposed to advance long-context capabilities: Architectural, e.g., to augment the embedding layer to cleverly extrapolate to unseen lengths (Vaswani et al., 2017; Press et al., 2021; Caciularu et al., 2022); and via data, e.g., to incorporate longer inputs and more challenging long-context scenarios into training (He et al., 2023; Chen et al., 2023). However, this challenging problem stubbornly remains in competitive models today (Liu et al., 2023; Bishop et al., 2023; Levy et al., 2024). In contrast to the above methods, DoubleDipper does not involve training or architectural changes.
In documenting and exploring LLM performance in long-context settings, many different benchmarks targeting it have been proposed, such as Scrolls and Zero-Scrolls (Shaham et al., 2022, 2023), Loogle (Li et al., 2023), LongBench (Bai et al., 2023), inter alia. We describe the most relevant evaluation benchmarks used in this work in Section 4.
In-Context Learning
The area of in-context learning (ICL) is a class of prompting techniques where demonstrations are added to the prompt in order to steer or improve model behavior (Min et al., 2022a). Typically, in-context learning involves hand-crafted demonstrations (Song et al., 2022), automatic retrieval of demonstrations from a larger set (Paranjape et al., 2023), or instructing the model to perform various tasks one after another in a pipeline (Gao et al., 2022). Recent improvements in long-context capabilities of LLMs have also had effect on improving the yield from in-context learning by simply using more short-length demonstrations (Agarwal et al., 2024).
While such methods are widely used for their effectiveness (Brown et al., 2020b; an Luo et al., 2024), they remain under-explored in settings of long-context. The reason is simple: If the context is already extremely long, adding additional demonstrations comparable in length to the input context will likely amplify the existing limitations of long context handling (Li et al., 2024). In this work, we tackle these challenges and present a novel ICL method for long contexts.
3 DoubleDipper
Figure 3: Example of DoubleDipper applied to the MuSique dataset. Given 20 passages as input, DoubleDipper randomly selects 3 passages (specifically passages 15, 5, 17) and automatically generates a question-pair for each one. As each QA is associated with its respective paragraph, we form the demonstrations to instruct the model to identify the relevant passage(s) and the correct answer.
Recall that our work focuses on the task of question answering (QA) with long input context comprising multiple paragraphs. In addition to the answer, we aim to identify the supporting paragraphs in order to provide attribution. Formally, given a long input text $D$ composed of $n$ paragraphs $D=\{p_{1},p_{2},...,p_{n}\}$ and a question $q$, the goal is to generate the answer $a$ and identify the set(s) of paragraphs that support the answer $S=\{s_{1},...,s_{k}\}$. The number of the supporting paragraphs is not known in advance and can be one or more.
We describe DoubleDipper, an efficient method for improving the performance of large language models (LLMs) when dealing with long contexts. The core principles of DoubleDipper involve: (1) recycling the input context to automatically generate few-shot examples, and (2) “teaching” the model via in-context learning (ICL) to explicitly pinpoint the supporting paragraphs before generating the answer.
Figure 3 illustrates DoubleDipper. Starting with the input paragraphs $D$, we initially select $k$ paragraphs at random (e.g., paragraphs $15$, $5$, and $17$, for $k:=3$). For each chosen paragraph, we prompt the model to formulate a question that pertains to the specific paragraph, accompanied by an appropriate answer (for further details on prompt specifications, refer to Appendix A). Each generated QA pair is directly associated with its origin paragraph, enabling us to assemble the following structured in-context demonstration, shown as the DoubleDipper block in Figure 3:
Question $:q_{i}$
Evidence $:p_{i}$
Answer $:a_{i}$
Here, $p_{i}$ indicates the index of the paragraph associated with the QA pair $(q_{i},a_{i})$. Given a test question $q$, we then form a QA prompt by concatenating the original input context $D$, the compiled demonstrations and $q$. The model first generates the one or more indices of the supporting paragraph(s), followed by the answer.
Unlike traditional few shot examples that instruct the model about a specific task, DoubleDipper aims to coach the model on how to “handle” the specific input context. This is achieved by guiding the model to explicitly localize relevant information before generating the answer. By randomly sampling multiple paragraphs from the input, DoubleDipper guarantees that the ICL demonstrations involve reading different parts of the context, allowing the model to better comprehend the input text. Beyond improving the performance of the QA task, instructing the model to provide the supporting paragraphs is valuable on its own as it offers transparency and substantially eases human evaluation (Slobodkin et al., 2024).
DoubleDipper offers several advantages. First, as each example in the demonstration consists only of a question, an answer and the ID of relevant passage, the number of added tokens due to the extra demonstrations is minimal, leading to a low additional cost and computation. Furthermore, by reusing the same context to generate demonstrations, our approach guarantees that all few shot examples are derived from the exact same domain as the input query (Rubin et al., 2022).
Finally, we observe that, although the QA pairs in the demonstration are confined to individual paragraphs, the actual query $q$ may require reasoning over multiple paragraphs (i.e., multi-hop QA). Surprisingly, LLMs can generalize from DoubleDipper local examples to these complex, global questions and successfully generate indices to multiple paragraphs (see Section 5).
4 Experiments
Datasets
Dataset # Instances Avg. # tokens
Lost-in-the-middle 2,500 2,815
FLenQA 1,500 3,225
HotpotQA 500 1,646
2Wiki 500 1,222
MuSiQue 500 2,549
Table 1: Evaluation datasets in our experiments. The average number of tokens is computed according to Gemma’s tokenization of the simple prompt.
We apply our method to various datasets, each presenting its own domain-specific challenges. We selected these datasets because the supporting paragraphs are also annotated. Overall our evaluation set include 5.5K instances, with statistics of each dataset given in Table 1.
The Lost-in-the-middle dataset (Liu et al., 2023) includes examples from NaturalQuestions-Open (Kwiatkowski et al., 2019; Lee et al., 2019). Each instance consists of twenty Wikipedia passages, with only one passage containing the answer to the query. The remaining passages are distractors that are lexically similar but do not contain the answer. To assess the robustness of large language models (LLMs) to the position of relevant information, Liu et al. (2023) evaluated cases where the relevant passage appeared in positions 1, 5, 10, 15, and 20. Following their methodology, we sampled 500 instances for each position, resulting in a total of 2,500 instances.
FLenQA (Levy et al., 2024) is a benchmark that includes simple questions with answers of either “True” or “False” based on two key sentences. FLenQA includes three subtasks. The first subtask is Monotone Relations (MonoRel), where each instance asks whether a transitive relation between two entities holds based on the context (e.g., "Is X younger than Y?" based on the sentences "X is younger than Z" and "Z is younger than Y"). The second subtask, People In Rooms (PIR), involves one key sentence indicating that a person is in a specific room and another key sentence describing a property of this room. The question asks whether the person is in a room with the described property. The final subtask is Simplified Rule Taker (SRT), based on RuleTaker (Clark et al., 2020). Each instance consists of a logical rule, two sentences each introducing a fact, and a question over the rule and facts. For each subtask, FLenQA includes contexts with varying lengths, from 50 to 3,000 tokens, by simply adding irrelevant text, demonstrating consistent performance degradation with increased input length. In our experiments, we sampled 250 instances for each subtask with input lengths of 2,000 and 3,000 tokens, leading to a total of 1,500 instances for FLenQA.
In addition, we evaluate our method on common multi-hop QA benchmarks. We sampled 500 instances from each of the following datasets: HotPotQA (Yang et al., 2018), 2Wiki (Ho et al., 2020), and MuSiQue (Trivedi et al., 2021). In all these datasets, the input text includes multiple passages, and models need to perform at least two steps of reasoning over different passages in order to answer the question.
Models
We apply our method to a variety of models, both commercial and open-source. The commercial models include Gemini-1.0-Ultra and Gemini-1.0-Pro (Team et al., 2023). The open-source models we tested are Llama-2-7b-chat, Llama-2-13b-chat, Llama-2-70b-chat (Touvron et al., 2023), Gemma-2b-it (v1.1), Gemma-7b-it (v1.1)(Team et al., 2024) and Mistral-7b-instruct (v0.2)(Jiang et al., 2023). Details about models’ size and context window are shown in Table 2.
For few-shot generation, we conducted two distinct experiments in order to analyze the effect of different question generation models. In the first experiment, dubbed DoubleDipper (Self), we employed the same LLM for both generating the demonstrations and answering to the query. In the second experiment, named DoubleDipper (PaLM 2), we generated a single set of few-shot examples using PaLM 2 (Anil et al., 2023) and then supplied these examples to various LLMs. This approach allowed us to assess the performance consistency across different models when provided with the same set of demonstrations.
Baselines
We compare DoubleDipper to the vanilla baseline, which takes as input the entire context $D$ and the query $q$ and generate only the answer $a$.
Size Context Window
Gemini v1.0 Pro 32k
Ultra 32k
Llama-2 7B 4k
13B 4k
70B 4k
Mistral 7B 8k
Gemma 2B 8k
7B 8k
Table 2: Models used with DoubleDipper.
Evaluation
We evaluate each dataset with the original evaluation metrics. Namely, we report Accuracy for Lost-in-the-middle (Liu et al., 2023) and FLenQA (Levy et al., 2024), and Token F1 for HotPotQA (Yang et al., 2018), 2Wiki (Ho et al., 2020) and MuSique (Trivedi et al., 2021).
In addition to the task’s accuracy, we also evaluate the performance of the identification of the supporting paragraph(s), by computing the F1 score on the predicted set of supporting passages compared to the ground truth (Yang et al., 2018; Ho et al., 2020; Trivedi et al., 2021).
Implementation Details
We randomly select three passages from the input (see Section 6 for an analysis of the number of self-generated demonstrations on the performance), each containing at least two sentences, and ask the model to generate five QA pairs per passage (see Appendix A for the exact prompt). We then randomly select a single QA pair for each passage to form the few shot demonstrations. For all experiments, including few-shot generation and question-answering, we use a temperature setting of 0.
Avg. Lost PIR MonoRel SRT HotPotQA 2Wiki MuSique
Gemini Pro 39.4 49.0 73.0 30.8 54.4 28.1 18.6 22.1
   + DoubleDipper (Self) 50.0 58.4 85.6 28.6 52.2 56.6 36.5 32.0
   + DoubleDipper (PaLM 2) 54.6 58.4 87.6 40.6 52.8 62.2 42.8 38.0
Gemini Ultra 53.9 53.0 81.2 30.8 66.0 61.6 42.6 42.3
   + DoubleDipper (Self) 53.1 57.2 73.4 20.6 54.6 68.9 51.5 45.6
   + DoubleDipper (PaLM 2) 54.7 56.8 78.6 25.0 56.6 69.1 52.6 44.2
Gemma-2b-it (v1.1) 35.9 24.0 53.4 49.8 49.6 34.6 28.1 11.6
   + DoubleDipper (Self) 35.5 25.5 66.0 54.2 49.2 24.7 18.6 10.5
   + DoubleDipper (PaLM 2) 41.2 23.6 75.8 57.8 49.4 38.2 28.1 15.4
Gemma-7b-it (v1.1) 27.2 7.1 52.2 49.8 48.6 12.3 16.7 3.4
   + DoubleDipper (Self) 47.6 45.8 92.6 74.2 45.4 40.4 22.4 12.4
   + DoubleDipper (PaLM 2) 51.3 44.0 94.8 74.8 45.2 49.6 32.1 18.4
Mistral-7b-instruct (v0.2) 45.8 61.3 70.6 68.0 49.0 36.3 21.7 13.9
   + DoubleDipper (Self) 48.9 60.4 96.2 84.6 42.6 28.3 17.8 12.6
   + DoubleDipper (PaLM 2) 52.8 58.9 95.6 82.8 44.4 43.1 26.4 18.4
Llama-2-7b-chat 31.6 42.0 46.8 39.0 37.4 32.7 15.1 8.2
   + DoubleDipper (Self) 33.3 45.6 64.2 37.2 35.4 23.3 17.8 9.2
   + DoubleDipper (PaLM 2) 38.5 43.9 67.6 36.8 35.2 39.2 30.1 16.4
Llama-2-13b-chat 32.4 51.2 54.2 40.6 43.8 18.2 10.8 7.9
   + DoubleDipper (Self) 34.9 46.9 66.0 44.6 38.4 22.7 19.3 8.4
   + DoubleDipper (PaLM 2) 38.3 47.6 71.0 37.8 36.4 32.8 32.2 10.4
Llama-2-70b-chat 45.3 61.1 66.6 76.4 50.2 33.8 17.7 11.0
   + DoubleDipper (Self) 46.1 57.2 85.6 76.6 44.8 27.8 20.3 10.7
   + DoubleDipper (PaLM 2) 51.6 54.2 88.2 78.0 47.0 44.5 34.2 15.0
Table 3: Results of the baseline, DoubleDipper (Self) and DoubleDipper (PaLM 2) on various QA datasets.
5 Results
Table 3 presents the QA performance of our two variants of DoubleDipper, namely DoubleDipper (Self) and DoubleDipper (PaLM 2) on our evaluation set. We report the performance of the supporting paragraphs prediction in Appendix B. On average, our straightforward method, which automatically generates few-shot examples from the provided input context, outperforms the baseline, often by a significant margin (e.g., +15.2 points for Gemini Pro, +7 points for Mistral Instruct, +6.3 points for Llama2 70B). Comparing the two variants, DoubleDipper (PaLM 2) generally yields better results than DoubleDipper (Self), indicating that higher-quality question-answer (QA) pairs in the few-shot examples enhance performance. Nonetheless, DoubleDipper (Self) already surpasses the baseline by a considerable margin on its own in most benchmarks. Notably, while DoubleDipper produces simple QAs answerable from a single paragraph, it demonstrates strong generalization across diverse QA formats (except SRT), including PIR and MonoRel, which involves True/False questions, and multi-hop QA datasets (HotPotQA, 2Wiki, and MuSique), where reasoning over multiple paragraphs is required.
Furthermore, following (Liu et al., 2023), Figure 4 shows the performance of the tested LLMs for both the baseline and DoubleDipper (Self) on our sample of the Lost-of-the-middle dataset, according to the position of the document that contains the answer. For Gemini Pro, Gemini Ultra, Gemma 2B, Gemma 7B, Llama 7B, the performance curve for DoubleDipper consistently surpasses the baseline across nearly all document positions. This reveals that beyond improving performance, DoubleDipper can make the model more robust to the position of the relevant document. For Mistral, Llama 2 13B and Llama 2 70B, Figure 4 shows that DoubleDipper performs on par or better than the baseline when the relevant document is in the middle of the input, but still underperforms the baseline at the beginning and the end. This variation can likely be attributed to the inherent biases of LLMs towards the beginning and end of inputs, while adding in context demonstrations mitigates this bias.
Figure 4: Performance (accuracy) of the models on our sample of the Lost-in-the-middle dataset (Liu et al., 2023) according to the position of the document that contains the answer.
6 Analysis
Few-shot vs. Zero-shot
Avg. Lost PIR MonoRel SRT HotPotQA 2Wiki MuSique
Gemini Pro Zero-shot 50.0 59.7 77.4 38.8 54.2 51.2 34.6 34.1
DoubleDipper 54.6 58.4 87.6 40.6 52.8 62.2 42.8 38.0
Gemini Ultra Zero-shot 56.5 56.2 76.8 34.8 62.2 68.3 52.4 44.9
DoubleDipper 54.7 56.8 78.6 25.0 56.6 69.1 52.6 44.2
Gemma-2b-it (v1.1) Zero-shot 3.6 6.4 4.2 5.4 0.0 3.3 4.9 0.7
DoubleDipper 41.2 23.6 75.8 57.8 49.4 38.2 28.1 15.4
Gemma-7b-it (v1.1) Zero-shot 7.5 2.6 36.0 5.4 0.0 3.9 3.3 1.0
DoubleDipper 51.3 44.0 94.8 74.8 45.2 49.6 32.1 18.4
Mistral-7b-instruct (v0.2) Zero-shot 41.9 59.0 77.4 67.4 51.8 13.7 12.3 11.8
DoubleDipper 52.8 58.9 95.6 82.8 44.4 43.1 26.4 18.4
Llama-2-7b-chat Zero-shot 22.4 29.2 45.0 29.2 33.4 9.8 7.9 2.4
DoubleDipper 38.5 43.9 67.6 36.8 35.2 39.2 30.1 16.4
Llama-2-13b-chat Zero-shot 14.0 8.0 43.0 3.0 37.0 5.2 0.3 1.2
DoubleDipper 38.0 36.7 71.4 36.6 50.0 27.4 27.4 16.8
Table 4: Comparison of DoubleDipper to prompting models in a zero-shot setting to provide explicitly the supporting paragraphs, then generate the answer.
In order to assess the effectiveness of the few shots in DoubleDipper, we prompt the model in a zero-shot setting to explicitly specify the ID of the relevant passage(s) before generating the answer. The results of the QA task are detailed in Table 4 and the performance in identifying supporting paragraphs is reported in Appendix B.
For all open-source models, this zero-shot prompting strategy leads to a significant decline in QA performance compared to the baseline (see Table 3), which merely prompts the model for the answer (e.g., -32.3 points for Gemma 2B, -9.2 points for Llama 7B, -4.2 points for Mistral). This drop is likely because retrieving relevant passages makes the task more complex compared to standard QA, as models are not typically trained to perform such “retrieval” tasks. This is further supported by the poor performance in supporting passages identification shown in Appendix B, where F1 scores are often close to zero.
Despite these challenges, DoubleDipper’s self-generated demonstrations systematically improve performance for all open-source models on both QA and supporting passages identification across all datasets. On average, Mistral’s QA performance improves by 10.9 points, and its identification of relevant paragraphs increased by 30.5 F1 score. This aligns with previous research suggesting that models can learn new tasks through in-context learning (ICL).
Comparison to traditional ICL
Avg. Lost PIR MonoRel SRT HotPotQA 2Wiki MuSique
Gemini Pro ICL 48.6 59.1 69.8 30.0 52.8 55.6 41.0 32.0
DoubleDipper 54.6 58.4 87.6 40.6 52.8 62.2 42.8 38.0
Gemma-2b-it (v1.1) ICL 21.6 0.2 28.2 26.4 13.0 39.7 39.1 4.3
DoubleDipper 41.2 23.6 75.8 57.8 49.4 38.2 28.1 15.4
Llama-2-7b-chat ICL 16.7 0 13.2 13.2 14.2 37.0 38.2 0.7
DoubleDipper 38.5 43.9 67.6 36.8 35.2 39.2 30.1 16.4
Mistral-7b-instruct ICL 52.1 68.5 79 67.8 39.2 49.4 36.7 24.1
DoubleDipper 52.8 58.9 95.6 82.8 44.4 43.1 26.4 18.4
Table 5: Comparison of traditional In-Context Learning (ICL) where each demonstration example comprises a full text, a question and an answer to DoubleDipper where the demonstrations contain only question-answer pairs, automatically generated on the same input text.
To analyze the performance gain of DoubleDipper compared to the traditional In-Context-Learning, we prompt a representative model from each family (Gemini Pro, Gemma 2B, Llama 7B, and Mistral) with few-shot examples from the same dataset, each example including its own context and question-answer pair.
We report the results in Table 5. For all models, DoubleDipper achieves a higher performance than the traditional ICL. Interestingly, for Gemma 2B and Llama 7B, the traditional ICL not only underperforms DoubleDipper but it also significantly hurts performance compared to the baseline (e.g., -14.3 for Gemma 2B and -14.9 for Llama 7b). This performance drop might be because each demonstration in the traditional ICL setup comprises a few thousand tokens, which exacerbates the models’ existing challenges with processing long-range dependencies.
As mentioned in Section 3, DoubleDipper promotes also efficiency by adding to the original prompt only a few extra tokens, leading to a significantly cheaper inference than the traditional ICL.
How many examples are needed?
In Table 6, we explore the impact of varying $k$, the number of self-generated few-shot examples in DoubleDipper to 3, 5, and 10. Our analysis reveals no significant differences in performance across these variations, while 3 self-generated examples are sufficient to improve performance. This finding is in line with previous work (Brown et al., 2020a; Min et al., 2022b).
$k=3$ $k=5$ $k=10$
Gemini Pro 54.6 54.1 54.7
Gemini Ultra 54.7 54.1 53.8
Gemma-2b-it (v1.1) 41.2 40.8 41.0
Gemma-7b-it (v1.1) 51.3 50.2 49.5
Mistral-7b-instruct (v0.2) 52.8 52.8 52.4
Llama-2-7b-chat 33.3 32.5 31.5
Llama-2-13b-chat 34.9 34.4 33.0
Llama-2-70b-chat 51.6 51.1 51.5
Table 6: Average performance on our evaluation set with various numbers of self-generated few shot demonstrations ($k$) in DoubleDipper.
DoubleDipper without identification of supporting paragraphs
To ablate the second principle in DoubleDipper, namely the explicit identification of the supporting paragraphs before generating the answer, we prompt Gemini Pro and Gemini Ultra with self-generated few shot examples that comprise only question-answer pairs (without instructing the model to retrieve the relevant passage(s)). Gemini Pro achieves on average 36.2 points (-18.4) and Gemini Ultra achieves 54.0 (-0.7 points). This indicates that instructing small models (e.g., Gemini Pro) with explicit prediction of supporting paragraphs achieves similar results to large models (Gemini Ultra).
Do few-shot examples really “instruct” the model to comprehend the text or provide clues to the query?
To answer this question, we prompt Gemini Pro and Gemini Ultra with only the self-generated demonstrations without the input context. As expected, removing the input context leads to a huge performance drop, -31.7 for Gemini Pro and -22.8 for Gemini Ultra compared to DoubleDipper. This confirms that DoubleDipper indeed “teaches” the model how to comprehend the input.
Qualitative Analysis
We manually analyze 50 prompts, with few-shot demonstrations produced by Gemini Ultra. Our review confirms that 93.5% of these self-generated QAs are correct, meaning that the question is meaningful and the answer could be found in the corresponding paragraph.
7 Conclusion
We develop DoubleDipper, a straightforward method for enhancing the performance of Question Answering with long context and providing attribution to the relevant paragraph(s) in the input. By recycling the input context to generate the few shot examples, each demonstration includes solely a question, an answer and a pointer to the relevant paragraph, without a separate context, thus effectively addressing the challenging of In-Context-Learning with long context. Experimental results show that our approach ourperforms both the vanilla LLM and the traditional ICL in various QA settings, including distractor passages in the input, True/False questions and multi-hop QA.
8 Limitations
One notable limitation of our approach is the extended inference time required for generating question-answer pairs. Future research could mitigate this issue by developing smaller, specialized models specifically tailored for QA generation.
Additionally, our evaluation set is constrained to instances that are solely in English and range between 1,000 to 4,000 tokens. Expanding the diversity of languages and token ranges could enhance the robustness and applicability of our findings.
Lastly, although we employ a strategy of randomly sampling $k$ paragraphs from the input to ensure the model engages with varied segments of the text, we did not optimize the selection of these paragraphs. Future work could explore more strategic methods for paragraph selection to potentially enhance the efficacy and relevance of the generated examples.
Appendix A Prompts
Figure 5 shows the zero-shot prompt we use for generating the question-answer pairs in DoubleDipper. For the QA prompts, we use the same instructions and prompt template as the original papers (Lost-in-the-middle and FLenQA) and add a simple line for the instructions in other multi-hop QA datasets: “Please answer the question based on the given passages below.”. For MuSique, since the dataset includes questions that are not answerable, we add the following sentence to the prompt: “If the question can’t be answered given the given passages, please write "unanswerable"”.
basicstyle=,xleftmargin=2.8em,xrightmargin=1.5em, showstringspaces=false, showspaces=false, showtabs=false, tabsize=2, breaklines=true, flexiblecolumns=true, escapeinside=<@@>, breakatwhitespace=true
Figure 5: Template prompt for the generation question-answer pairs.
Appendix B Identification of Supporting Passages
Table 7 presents the F1 results of the tested models for the supporting relevant passages identification. Without any demonstration (zero-shot), all open source models achieve a poor performance, while DoubleDipper significantly improves performance (+37.8 F1 for Llama2 70B, +38 F1 for Gemma 7B, +30.5 F1 for Mistral).
Avg. Lost PIR MonoRel SRT HotPotQA 2Wiki MuSique
Gemini Pro Baseline - - - - - - - -
Zero-shot 54.5 58.2 47.1 75.9 46.6 56.3 58.3 39.4
DoubleDipper (Self) 50.0 59.7 77.4 38.8 54.2 51.2 34.6 34.1
DoubleDipper (PaLM 2) 50.6 60.1 47.3 68.0 21.2 58.8 57.9 41.0
Gemini Ultra Baseline - - - - - - - -
Zero-shot 40.4 57.4 30.3 39.4 17.0 52.4 43.7 42.9
DoubleDipper (Self) 37.5 58.6 25.5 21.5 11.1 55.4 46.4 44.3
DoubleDipper (PaLM 2) 38.4 58.0 28.0 24.6 12.0 55.3 47.0 43.7
Gemma-2b-it (v1.1) Baseline - - - - - - - -
Zero-shot 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
DoubleDipper (Self) 23.3 10.4 43.1 37.5 13.5 20.1 13.0 25.3
DoubleDipper (PaLM 2) 25.6 9.1 48.7 48.7 15 18.7 14 25.3
Gemma-7b-it (v1.1) Baseline - - - - - - - -
Zero-shot 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
DoubleDipper (Self) 38.0 39.7 44.8 59.8 7.1 41.8 36.5 36.1
DoubleDipper (PaLM 2) 36.9 39.2 45.2 58.6 7.3 40.1 32.8 35
Mistral-7b-instruct (v0.2) Baseline - - - - - - - -
Zero-shot 11.5 6.7 0.7 2.9 0.0 17.2 18.2 34.5
DoubleDipper (Self) 42.0 48.2 66.2 60.6 5.8 41.7 29.3 42.2
DoubleDipper (PaLM 2) 41.5 48.3 70.0 58.6 6.8 40.0 26.0 40.6
Llama-2-7b-chat Baseline - - - - - - - -
Zero-shot 0.8 0.2 0.0 0.0 3.2 0.0 0.0 2.1
DoubleDipper (Self) 19.4 17.5 27.7 20.2 5.7 21.7 19.7 23.3
DoubleDipper (PaLM 2) 19.6 15.8 29.3 21.6 5.7 22.2 20.7 22.0
Llama-2-13b-chat Baseline - - - - - - - -
Zero-shot 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
DoubleDipper (Self) 28.8 16.5 51.0 32.0 18.4 29.6 22.2 31.7
DoubleDipper (PaLM 2) 30.3 13.4 51.6 30.5 15.6 36.4 27.3 36.9
Table 7: Performance (F1) of supporting paragraph(s) prediction. The baseline does not predict supporting paragraphs. DoubleDipper provides a significant performance boost for all open source models, whose performance is close to 0 (except from Mistral) in the zero shot setting.

### v2

@@
 We further enhance each demonstration by instructing the model to explicitly identify the relevant paragraphs before the answer, which improves performance while providing fine-grained attribution to the answer source.
-We apply our method on multiple LLMs and obtain substantial improvements on various QA datasets with long context, especially when the answer lies within the middle of the context.
+We apply our method on multiple LLMs and obtain substantial improvements (+23% on average across models) on various QA datasets with long context, especially when the answer lies within the middle of the context.
 Surprisingly, despite introducing only single-hop ICL examples, LLMs also successfully generalize to multi-hop long-context QA using our approach.
@@
 We applied DoubleDipper to a variety of LLMs, both commercial (Gemini Pro and Gemini Ultra; Team et al., 2023) and open-source (Llama (Touvron et al., 2023), Mistral (Jiang et al., 2023) and Gemma (Team et al., 2024)), and evaluate it on various QA datasets with long inputs, including common multi-hop QA datasets.
-Our experiments demonstrate that with only 3 self-generated few-shot examples, DoubleDipper consistently outperforms the baseline on our evaluation set.
+Our experiments demonstrate that with only 3 self-generated few-shot examples, DoubleDipper consistently outperforms the baseline on our evaluation set by 23% on average across models.
 In addition, for some models, DoubleDipper enhances the robustness to the position of the relevant information within the text.
-Interestingly, while our few-shot examples focus on single-paragraph answers, DoubleDipper generalizes well to multi-hop QAs where the answer requires information from multiple passages.
+Interestingly, while our few-shot examples focus on single-paragraph answers, DoubleDipper generalizes well to multi-hop QAs and where the answer requires information from multiple passages.
 2 Background
@@
 We report the performance of the supporting paragraphs prediction in Appendix B.
-On average, our straightforward method, which automatically generates few-shot examples from the provided input context, outperforms the baseline, often by a significant margin (e.g., +15.2 points for Gemini Pro, +7 points for Mistral Instruct, +6.3 points for Llama2 70B).
-Comparing the two variants, DoubleDipper (PaLM 2) generally yields better results than DoubleDipper (Self), indicating that higher-quality question-answer (QA) pairs in the few-shot examples enhance performance.
-Nonetheless, DoubleDipper (Self) already surpasses the baseline by a considerable margin on its own in most benchmarks.
-Notably, while DoubleDipper produces simple QAs answerable from a single paragraph, it demonstrates strong generalization across diverse QA formats (except SRT), including PIR and MonoRel, which involves True/False questions, and multi-hop QA datasets (HotPotQA, 2Wiki, and MuSique), where reasoning over multiple paragraphs is required.
+On average across models and datasets, our straightforward method, which automatically generates few-shot examples from the provided input context, yields an improvement of 12% for DoubleDipper (Self) and 23% for DoubleDipper (PaLM 2), often by a significant margin (e.g., +15.2 points for Gemini Pro, +7 points for Mistral Instruct, +6.3 points for Llama2 70B).
+Notably, while DoubleDipper produces simple QAs answerable from a single paragraph, it always surpasses the baseline in multi-hop QA datasets (HotPotQA, 2Wiki and MuSique).
 Furthermore, following (Liu et al., 2023), Figure 4 shows the performance of the tested LLMs for both the baseline and DoubleDipper (Self) on our sample of the Lost-of-the-middle dataset, according to the position of the document that contains the answer.

### v3

@@
 We further enhance each demonstration by instructing the model to explicitly identify the relevant paragraphs before the answer, which improves performance while providing fine-grained attribution to the answer source.
-We apply our method on multiple LLMs and obtain substantial improvements (+23% on average across models) on various QA datasets with long context, especially when the answer lies within the middle of the context.
+We apply our method on multiple LLMs and obtain substantial improvements (+16 absolute points on average across models) on various QA datasets with long context, especially when the answer lies within the middle of the context.
 Surprisingly, despite introducing only single-hop ICL examples, LLMs also successfully generalize to multi-hop long-context QA using our approach.
@@
 1 Introduction
-Figure 1: Performance of Gemini Pro (v1.0) on a sample of the Lost-in-the-middle dataset (Liu et al., 2023).
+Figure 1: Performance of Gemini Flash (v1.5) on a sample of the Lost-in-the-middle dataset (Liu et al., 2023).
 The X-axis is the position of the relevant passage in the context.
 The baseline (blue line) displays a U-shaped curve, performing well only when the relevant passage is at the beginning or end of the input.
-The oracle (green line) shows significant performance gain when the relevant passage ID is provided in the prompt, showing that the main bottleneck is the identification of supporting evidence(s).
+The oracle (green line) shows significant performance gain when the relevant passage ID is provided in the prompt, showing that the identification of supporting evidence(s) is a major challenge.
 DoubleDipper (our method, orange line) flattens this U-shaped trend.
@@
 It has been shown that LLMs struggle when the relevant information is buried in the middle of the context (Liu et al., 2023) or obscured by numerous irrelevant details (Levy et al., 2024).
-Our analysis (illustrated in Figure 1) identifies the identification of relevant information as a primary performance bottleneck of current models in long contexts.
+Our analysis (illustrated in Figure 1) identifies the identification of relevant information as a major performance bottleneck of current models in long contexts.
 In this work, we introduce a novel method to enhance the QA performance of LLMs in long input setups (to allow direct comparisons across a wide swath of models, we limit "long context" here to 1-3k tokens).
@@
 Finally, DoubleDipper generates answers with attribution to relevant paragraphs, improving the model’s lookup ability and offering transparency, which substantially simplifies human evaluation (Slobodkin et al., 2024).
-We applied DoubleDipper to a variety of LLMs, both commercial (Gemini Pro and Gemini Ultra; Team et al., 2023) and open-source (Llama (Touvron et al., 2023), Mistral (Jiang et al., 2023) and Gemma (Team et al., 2024)), and evaluate it on various QA datasets with long inputs, including common multi-hop QA datasets.
-Our experiments demonstrate that with only 3 self-generated few-shot examples, DoubleDipper consistently outperforms the baseline on our evaluation set by 23% on average across models.
+We applied DoubleDipper to a variety of LLMs, both commercial (Gemini Pro, Nano and Flash; Reid et al., 2024) and open-source (Llama 3.1 (Dubey et al., 2024), Mistral (Jiang et al., 2023), Mixtral (Jiang et al., 2024) and Gemma (Riviere et al., 2024)), and evaluate it on various QA datasets with long inputs, including common multi-hop QA datasets.
+Our experiments demonstrate that with only 3 self-generated few-shot examples, DoubleDipper consistently outperforms the baseline on our evaluation set by 16 absolute points on average across models.
 In addition, for some models, DoubleDipper enhances the robustness to the position of the relevant information within the text.
@@
 LLMs have been well-documented to struggle when input length grows (An et al., 2023), and especially so when it exceeds input lengths seen during training (Anil et al., 2022).
-Various methods have been proposed to advance long-context capabilities: Architectural, e.g., to augment the embedding layer to cleverly extrapolate to unseen lengths (Vaswani et al., 2017; Press et al., 2021; Caciularu et al., 2022); and via data, e.g., to incorporate longer inputs and more challenging long-context scenarios into training (He et al., 2023; Chen et al., 2023).
+Various methods have been proposed to advance long-context capabilities: Architectural, e.g., to augment the embedding layer to cleverly extrapolate to unseen lengths (Vaswani et al., 2017; Press et al., 2021; Caciularu et al., 2022; Hsieh et al., 2024); and via data, e.g., to incorporate longer inputs and more challenging long-context scenarios into training (He et al., 2023; Chen et al., 2023).
 However, this challenging problem stubbornly remains in competitive models today (Liu et al., 2023; Bishop et al., 2023; Levy et al., 2024).
@@
 In documenting and exploring LLM performance in long-context settings, many different benchmarks targeting it have been proposed, such as Scrolls and Zero-Scrolls (Shaham et al., 2022, 2023), Loogle (Li et al., 2023), LongBench (Bai et al., 2023), inter alia.
+The problem of designing informative and reliable benchmarks in long-context is an an active, ever-changing area of research (Goldman et al., 2024; Yen et al., 2024).
 We describe the most relevant evaluation benchmarks used in this work in Section 1.
@@
 While such methods are widely used for their effectiveness (Brown et al., 2020b; an Luo et al., 2024), they remain under-explored in settings of long-context.
-The reason is simple: If the context is already extremely long, adding additional demonstrations comparable in length to the input context will likely amplify the existing limitations of long context handling (Li et al., 2024).
-In this work, we tackle these challenges and present a novel ICL method for long contexts.
+The reason is simple: If the context is already extremely long, adding additional demonstrations comparable in length to the input context will likely amplify the existing limitations of long context handling (Li et al., 2024c).
+More related to our work, a few recent studies propose to prompt LLMs for automatically generating in-context demonstrations for various short context tasks (Kim et al., 2022; Yasunaga et al., 2024; Li et al., 2024a, b).
+However, applying those methods to long contexts is not straightforward because LLMs would need to generate long demonstrations.
+In this work, we tackle these challenges and present a novel ICL method that recycles the given long context and further instructs models to identify the relevant information before generating the answer.
 3 DoubleDipper
@@
 This is achieved by guiding the model to explicitly localize relevant information before generating the answer.
-By randomly sampling multiple paragraphs from the input, DoubleDipper guarantees that the ICL demonstrations involve reading different parts of the context, allowing the model to better comprehend the input text.
+Also, by randomly sampling multiple paragraphs from the input, DoubleDipper guarantees that the ICL demonstrations involve reading different parts of the context, allowing the model to better comprehend the input text.
 Beyond improving the performance of the QA task, instructing the model to provide the supporting paragraphs is valuable on its own as it offers transparency and substantially eases human evaluation (Slobodkin et al., 2024).
 DoubleDipper offers several advantages.
-First, as each example in the demonstration consists only of a question, an answer and the ID of relevant passage, the number of added tokens due to the extra demonstrations is minimal, leading to a low additional cost and computation.
+First, as each example in the demonstration consists only of a question, an answer and the ID of relevant passage, the number of added tokens due to the extra demonstrations is minimal, leading to a low additional cost and computation compared to the traditional In-Context-Learning.
 Furthermore, by reusing the same context to generate demonstrations, our approach guarantees that all few shot examples are derived from the exact same domain as the input query (Rubin et al., 2022).
@@
 4 Experiments
+Size Context Window
+Gemini v1.5 Nano 32k
+Flash 32k
+Pro 32k
+Llama 3.1 7B 128k
+Mistral v0.3 7B 32k
+Mixtral v0.1 7x8B 32k
+Gemma 2 2B 8k
+9B 8k
+Table 1: Models used with DoubleDipper.
 Datasets
@@
 We selected these datasets because the supporting paragraphs are also annotated.
-Overall our evaluation set include 5.5K instances, with statistics of each dataset given in Table 1.
+Overall our evaluation set includes 5.5K instances, with statistics of each dataset given in Table 2.
 The Lost-in-the-middle dataset (Liu et al., 2023) includes examples from NaturalQuestions-Open (Kwiatkowski et al., 2019; Lee et al., 2019).
@@
 In all these datasets, the input text includes multiple passages, and models need to perform at least two steps of reasoning over different passages in order to answer the question.
+Avg. 2Wiki MonoRel PIR SRT HotPotQA Lost MuSique
+Gemini Pro (vanilla) 60.5 24.9 95.0 97.6 64.4 46.7 71.6 23.1
+   Zero-shot + Evidence Retrieval 62.3 32.5 94.6 95.8 62.4 49.6 74.8 26.5
+   DoubleDipper 70.4 46.8 97.4 99.0 79.6 60.9 72.4 36.4
+Gemini Flash (vanilla) 42.9 10.2 70.0 86.0 57.6 10.0 59.5 7.3
+   Zero-shot + Evidence Retrieval 58.2 30.2 78.8 90.6 65.0 44.9 67.4 30.2
+   DoubleDipper 66.1 48.0 85.8 95.0 68.6 60.6 65.0 39.7
+Gemini Nano (vanilla) 41.6 10.8 72.2 66.8 55.4 21.3 59.6 5.2
+   Zero-shot + Evidence Retrieval 56.5 32.0 82.4 82.4 56.4 56.7 60.5 25.2
+   DoubleDipper 62.1 40.6 86.6 95.4 56.2 65.1 60.5 30.4
+Gemma 2 2B (v2) (vanilla) 38.6 8.9 71.8 68.6 51.2 13.3 49.6 6.5
+   Zero-shot + Evidence Retrieval 42.0 22.3 66.8 70.6 40.2 30.6 47.6 16.2
+   DoubleDipper 49.5 23.7 85.8 81.6 50.0 39.9 46.7 18.8
+Gemma 2 9B (v2) (vanilla) 44.0 11.4 74.8 81.8 55.6 13.8 61.0 9.3
+   Zero-shot + Evidence Retrieval 58.7 38.6 82.0 83.4 59.4 56.1 64.4 26.8
+   DoubleDipper 61.2 41.7 84.0 95.0 51.4 61.2 61.8 33.3
+Llama 3.1 8B (vanilla) 37.2 11.8 56.2 52.2 48.6 20.7 63.5 7.4
+   Zero-shot + Evidence Retrieval 53.1 42.2 71.8 65.4 50.8 55.3 62.0 24.5
+   DoubleDipper 59.9 38.7 91.2 90.6 51.0 59.3 58.1 30.1
+Mistral 7B (v0.3) (vanilla) 37.4 14.1 59.2 57.4 50.2 15.8 60.8 4.6
+   Zero-shot + Evidence Retrieval 44.0 23.8 66.2 62.4 49.6 34.1 58.6 13.6
+   DoubleDipper 51.0 28.6 68.4 88.8 50.6 43.4 60.7 16.7
+Mixtral 7x8B (v0.1) (vanilla) 42.6 13.7 73.0 66.2 51.0 18.2 67.7 8.4
+   Zero-shot + Evidence Retrieval 47.4 18.8 81.8 73.6 50.6 26.3 67.9 13.1
+   DoubleDipper 52.2 22.3 91.8 86.0 47.8 35.1 66.6 16.0
+Table 3: Accuracy of the QA task for the vanilla baseline (prompting the model to only answer the question), Zero-shot + Evidence Retrieval (prompting the model to explicitly identify the relevant passage(s) before generating the answer) and DoubleDipper with 3 demonstrations generated by Gemma 2 2B.
 Models
-We apply our method to a variety of models, both commercial and open-source.
-The commercial models include Gemini-1.0-Ultra and Gemini-1.0-Pro (Team et al., 2023).
-The open-source models we tested are Llama-2-7b-chat, Llama-2-13b-chat, Llama-2-70b-chat (Touvron et al., 2023), Gemma-2b-it (v1.1), Gemma-7b-it (v1.1)(Team et al., 2024) and Mistral-7b-instruct (v0.2)(Jiang et al., 2023).
+We apply DoubleDipper to a variety of models, both commercial and open-source.
+The commercial models include Gemini 1.5 Pro, Gemini 1.5 Flash and Gemini 1.0 Nano (Reid et al., 2024).
+The open-source models we tested are Llama 3.1 8B (Dubey et al., 2024), Gemma 2B (v2) and Gemma 9B (v2) (Riviere et al., 2024), Mistral-7B-Instruct (v0.2) (Jiang et al., 2023) and Mixtral-8x7B-Instruct (v0.1) (Jiang et al., 2024).
 Details about models’ size and context window are shown in Table 1.
-For few-shot generation, we conducted two distinct experiments in order to analyze the effect of different question generation models.
-In the first experiment, dubbed DoubleDipper (Self), we employed the same LLM for both generating the demonstrations and answering to the query.
-In the second experiment, named DoubleDipper (PaLM 2), we generated a single set of few-shot examples using PaLM 2 (Anil et al., 2023) and then supplied these examples to various LLMs.
-This approach allowed us to assess the performance consistency across different models when provided with the same set of demonstrations.
+Few-shot generation in DoubleDipper is an auxiliary task and should ideally run in an efficient time without requiring heavy resources.
+Therefore, in our main experiments, we employ Gemma 2B to generate the demonstrations at it is the smallest and most efficient model used in our experiments.
+See Section 6 for an ablation analysis of the effect of the chosen model for generating the demonstrations.
 Baselines
-We compare DoubleDipper to the vanilla baseline, which takes as input the entire context $D$ and the query $q$ and generate only the answer $a$.
-Size Context Window
-Gemini v1.0 Pro 32k
-Ultra 32k
-Llama-2 7B 4k
-13B 4k
-70B 4k
-Mistral 7B 8k
-Gemma 2B 8k
-7B 8k
-Table 2: Models used with DoubleDipper.
+We compare DoubleDipper to the vanilla baseline, which takes as input the entire context $D$ and the query $q$ and generates only the answer $a$.
+We also compare to Zero-shot + Evidence Retrieval, which prompts the model in a zero-shot setting to identify the relevant passage(s) before generating the answer.
 Evaluation
@@
 Implementation Details
-We randomly select three passages from the input (see Section 6 for an analysis of the number of self-generated demonstrations on the performance), each containing at least two sentences, and ask the model to generate five QA pairs per passage (see Appendix A for the exact prompt).
-We then randomly select a single QA pair for each passage to form the few shot demonstrations.
+We randomly select three passages from the input (see Section 6 for an analysis of the number of self-generated demonstrations on the performance), each containing at least two sentences, and ask the model to generate a single QA pair for each passage (see Appendix A for the exact prompt).
 For all experiments, including few-shot generation and question-answering, we use a temperature setting of 0.
-Avg. Lost PIR MonoRel SRT HotPotQA 2Wiki MuSique
-Gemini Pro 39.4 49.0 73.0 30.8 54.4 28.1 18.6 22.1
-   + DoubleDipper (Self) 50.0 58.4 85.6 28.6 52.2 56.6 36.5 32.0
-   + DoubleDipper (PaLM 2) 54.6 58.4 87.6 40.6 52.8 62.2 42.8 38.0
-Gemini Ultra 53.9 53.0 81.2 30.8 66.0 61.6 42.6 42.3
-   + DoubleDipper (Self) 53.1 57.2 73.4 20.6 54.6 68.9 51.5 45.6
-   + DoubleDipper (PaLM 2) 54.7 56.8 78.6 25.0 56.6 69.1 52.6 44.2
-Gemma-2b-it (v1.1) 35.9 24.0 53.4 49.8 49.6 34.6 28.1 11.6
-   + DoubleDipper (Self) 35.5 25.5 66.0 54.2 49.2 24.7 18.6 10.5
-   + DoubleDipper (PaLM 2) 41.2 23.6 75.8 57.8 49.4 38.2 28.1 15.4
-Gemma-7b-it (v1.1) 27.2 7.1 52.2 49.8 48.6 12.3 16.7 3.4
-   + DoubleDipper (Self) 47.6 45.8 92.6 74.2 45.4 40.4 22.4 12.4
-   + DoubleDipper (PaLM 2) 51.3 44.0 94.8 74.8 45.2 49.6 32.1 18.4
-Mistral-7b-instruct (v0.2) 45.8 61.3 70.6 68.0 49.0 36.3 21.7 13.9
-   + DoubleDipper (Self) 48.9 60.4 96.2 84.6 42.6 28.3 17.8 12.6
-   + DoubleDipper (PaLM 2) 52.8 58.9 95.6 82.8 44.4 43.1 26.4 18.4
-Llama-2-7b-chat 31.6 42.0 46.8 39.0 37.4 32.7 15.1 8.2
-   + DoubleDipper (Self) 33.3 45.6 64.2 37.2 35.4 23.3 17.8 9.2
-   + DoubleDipper (PaLM 2) 38.5 43.9 67.6 36.8 35.2 39.2 30.1 16.4
-Llama-2-13b-chat 32.4 51.2 54.2 40.6 43.8 18.2 10.8 7.9
-   + DoubleDipper (Self) 34.9 46.9 66.0 44.6 38.4 22.7 19.3 8.4
-   + DoubleDipper (PaLM 2) 38.3 47.6 71.0 37.8 36.4 32.8 32.2 10.4
-Llama-2-70b-chat 45.3 61.1 66.6 76.4 50.2 33.8 17.7 11.0
-   + DoubleDipper (Self) 46.1 57.2 85.6 76.6 44.8 27.8 20.3 10.7
-   + DoubleDipper (PaLM 2) 51.6 54.2 88.2 78.0 47.0 44.5 34.2 15.0
-Table 3: Results of the baseline, DoubleDipper (Self) and DoubleDipper (PaLM 2) on various QA datasets.
 5 Results
-Table 3 presents the QA performance of our two variants of DoubleDipper, namely DoubleDipper (Self) and DoubleDipper (PaLM 2) on our evaluation set.
-We report the performance of the supporting paragraphs prediction in Appendix B.
-On average across models and datasets, our straightforward method, which automatically generates few-shot examples from the provided input context, yields an improvement of 12% for DoubleDipper (Self) and 23% for DoubleDipper (PaLM 2), often by a significant margin (e.g., +15.2 points for Gemini Pro, +7 points for Mistral Instruct, +6.3 points for Llama2 70B).
+Result 1: DoubleDipper offers a substantial performance boost.
+Table 3 presents the QA performance of the baseline, Zero-shot + Evidence Retrieval and DoubleDipper on our evaluation set.
+The results first show that prompting models to explicitly identify the relevant paragraphs before generating the answer (Zero-shot + Evidence Retrieval) leads to a performance improvement of 9.7 points on average across models over the vanilla baseline.
+DoubleDipper offers an additional substantial boost of 6.3 points for all models on average, culminating in a overall improvement of 16 absolute points over the vanilla baseline.
 Notably, while DoubleDipper produces simple QAs answerable from a single paragraph, it always surpasses the baseline in multi-hop QA datasets (HotPotQA, 2Wiki and MuSique).
-Furthermore, following (Liu et al., 2023), Figure 4 shows the performance of the tested LLMs for both the baseline and DoubleDipper (Self) on our sample of the Lost-of-the-middle dataset, according to the position of the document that contains the answer.
-For Gemini Pro, Gemini Ultra, Gemma 2B, Gemma 7B, Llama 7B, the performance curve for DoubleDipper consistently surpasses the baseline across nearly all document positions.
+Likewise, DoubleDipper outperforms the baseline also on the FLenQA datasets (PIR, MonoRel and SRT), which involve synthetic True/False questions although the demonstrations in DoubleDipper are typically simple factoid questions.
+Avg. 2Wiki MonoRel PIR SRT HotPotQA Lost MuSique
+Gemini Pro Zero-shot + Evidence Retrieval 83.7 96.7 97.7 97.5 62.8 92.1 63.6 75.3
+DoubleDipper 86.3 94.4 99.8 97.1 80.9 90.0 66.4 75.4
+Gemini Flash Zero-shot + Evidence Retrieval 75.7 82.3 90.5 72.6 70.4 80.9 67.3 65.9
+DoubleDipper 79.1 83.7 98.3 80.2 71.2 84.6 66.1 69.5
+Gemini Nano Zero-shot + Evidence Retrieval 68.1 76.3 77.6 71.6 66.0 74.1 55.8 55.2
+DoubleDipper 69.9 76.6 86.6 74.5 59.8 76.0 59.2 56.7
+Gemma 2B Zero-shot + Evidence Retrieval 39.1 57.6 49.4 44.3 18.9 53.1 14.0 36.4
+DoubleDipper 45.6 56.5 60.7 57.5 13.1 58.1 33.1 39.9
+Gemma 9B Zero-shot + Evidence Retrieval 61.9 76.7 69.3 60.3 43.2 76.5 51.5 55.9
+DoubleDipper 57.0 74.2 52.5 59.0 23.1 78.4 55.3 56.8
+Llama 3.1 8B Zero-shot + Evidence Retrieval 61.7 53.2 86.5 66.9 67.8 63.2 41.1 53.5
+DoubleDipper 61.4 68.9 71.4 54.9 52.8 73.7 53.0 54.8
+Mistral 7B (v0.3) Zero-shot + Evidence Retrieval 46.6 62.4 49.8 46.2 17.5 64.0 43.4 43.0
+DoubleDipper 43.8 63.4 33.8 42.5 4.2 66.7 55.6 40.2
+Mixtral 7x8B v(0.1) Zero-shot + Evidence Retrieval 60.0 72.4 69.4 64.6 43.4 76.9 40.6 52.4
+DoubleDipper 58.9 70.4 81.6 63.6 18.3 75.0 50.4 53.2
+Table 4: Performance (F1) of supporting paragraph(s) prediction.
+Figure 4: Performance (accuracy) of Gemma 2 9B, Mixtral 8x7B and Gemini Flash with and without DoubleDipper on our sample of the Lost-in-the-middle dataset (Liu et al., 2023) according to the position of the document that contains the answer.
+Result 2: Learning to retrieve the evidence(s) with DoubleDipper is more effective in commercial models than in open source.
+Table 4 presents the performance of the supporting paragraphs prediction for the Zero-shot + Evidence Retrieval and DoubleDipper on our evaluation set.
+For all commercial models and Gemma 2B, DoubleDipper predicts better the supporting paragraphs than in the zero-shot setting (+2.6 F1 for Gemini Pro, +3.4 F1 for Gemini Flash, +1.8 F1 for Gemini Nano and +6.5 F1 for Gemma 2B).
+On the other hand, DoubleDipper slightly hurts the performance of common open source models (e.g., -2.8 F1 for Mistral and -0.3 F1 for Llama 3.1).
+This difference is because DoubleDipper’s demonstrations are based on a single paragraph, where commercial models can generalize better and predict multiple evidences.
+Indeed, Gemini Pro predicts on average 2 evidences for the datasets 2Wiki, MuSique, PIR, MonoRel, SRT and HotPotQA whereas Gemma 9B predicts 1.2 evidences for each instance.
+In fact, the only dataset with a single evidence is “Lost” and almost all models highly benefit from DoubleDipper (e.g, +11.9 F1 for Llama 3.1, +12.2 F1 for Mistral 7B and +9.8 for Mixtral 7x8B).
+Result 3: DoubleDipper makes models more robust to the position of relevant information.
+Following Liu et al.
+(2023), Figure 4 shows the performance of Gemma 2 9B, Mixtral 8x7B and Gemini Flash for both the baseline and DoubleDipper on our sample of the Lost-of-the-middle dataset, according to the position of the document that contains the answer.
+See Appendix B for the performance curve of the other tested models, which show similar trends.
+Overall, the performance curve for DoubleDipper consistently surpasses the baseline when the relevant information appears “in the middle” and sometimes also at the beginning and/or the end (e.g., Gemini Flash).
+This variation can likely be attributed to the inherent biases of LLMs towards the beginning and end of inputs, while adding in context demonstrations mitigates this bias.
 This reveals that beyond improving performance, DoubleDipper can make the model more robust to the position of the relevant document.
-For Mistral, Llama 2 13B and Llama 2 70B, Figure 4 shows that DoubleDipper performs on par or better than the baseline when the relevant document is in the middle of the input, but still underperforms the baseline at the beginning and the end.
-This variation can likely be attributed to the inherent biases of LLMs towards the beginning and end of inputs, while adding in context demonstrations mitigates this bias.
-Figure 4: Performance (accuracy) of the models on our sample of the Lost-in-the-middle dataset (Liu et al., 2023) according to the position of the document that contains the answer.
-6 Analysis
-Few-shot vs. Zero-shot
-Avg. Lost PIR MonoRel SRT HotPotQA 2Wiki MuSique
-Gemini Pro Zero-shot 50.0 59.7 77.4 38.8 54.2 51.2 34.6 34.1
-DoubleDipper 54.6 58.4 87.6 40.6 52.8 62.2 42.8 38.0
-Gemini Ultra Zero-shot 56.5 56.2 76.8 34.8 62.2 68.3 52.4 44.9
-DoubleDipper 54.7 56.8 78.6 25.0 56.6 69.1 52.6 44.2
-Gemma-2b-it (v1.1) Zero-shot 3.6 6.4 4.2 5.4 0.0 3.3 4.9 0.7
-DoubleDipper 41.2 23.6 75.8 57.8 49.4 38.2 28.1 15.4
-Gemma-7b-it (v1.1) Zero-shot 7.5 2.6 36.0 5.4 0.0 3.9 3.3 1.0
-DoubleDipper 51.3 44.0 94.8 74.8 45.2 49.6 32.1 18.4
-Mistral-7b-instruct (v0.2) Zero-shot 41.9 59.0 77.4 67.4 51.8 13.7 12.3 11.8
-DoubleDipper 52.8 58.9 95.6 82.8 44.4 43.1 26.4 18.4
-Llama-2-7b-chat Zero-shot 22.4 29.2 45.0 29.2 33.4 9.8 7.9 2.4
-DoubleDipper 38.5 43.9 67.6 36.8 35.2 39.2 30.1 16.4
-Llama-2-13b-chat Zero-shot 14.0 8.0 43.0 3.0 37.0 5.2 0.3 1.2
-DoubleDipper 38.0 36.7 71.4 36.6 50.0 27.4 27.4 16.8
-Table 4: Comparison of DoubleDipper to prompting models in a zero-shot setting to provide explicitly the supporting paragraphs, then generate the answer.
-In order to assess the effectiveness of the few shots in DoubleDipper, we prompt the model in a zero-shot setting to explicitly specify the ID of the relevant passage(s) before generating the answer.
-The results of the QA task are detailed in Table 4 and the performance in identifying supporting paragraphs is reported in Appendix B.
-For all open-source models, this zero-shot prompting strategy leads to a significant decline in QA performance compared to the baseline (see Table 3), which merely prompts the model for the answer (e.g., -32.3 points for Gemma 2B, -9.2 points for Llama 7B, -4.2 points for Mistral).
-This drop is likely because retrieving relevant passages makes the task more complex compared to standard QA, as models are not typically trained to perform such “retrieval” tasks.
-This is further supported by the poor performance in supporting passages identification shown in Appendix B, where F1 scores are often close to zero.
-Despite these challenges, DoubleDipper’s self-generated demonstrations systematically improve performance for all open-source models on both QA and supporting passages identification across all datasets.
-On average, Mistral’s QA performance improves by 10.9 points, and its identification of relevant paragraphs increased by 30.5 F1 score.
-This aligns with previous research suggesting that models can learn new tasks through in-context learning (ICL).
-Comparison to traditional ICL
-Avg. Lost PIR MonoRel SRT HotPotQA 2Wiki MuSique
-Gemini Pro ICL 48.6 59.1 69.8 30.0 52.8 55.6 41.0 32.0
-DoubleDipper 54.6 58.4 87.6 40.6 52.8 62.2 42.8 38.0
-Gemma-2b-it (v1.1) ICL 21.6 0.2 28.2 26.4 13.0 39.7 39.1 4.3
-DoubleDipper 41.2 23.6 75.8 57.8 49.4 38.2 28.1 15.4
-Llama-2-7b-chat ICL 16.7 0 13.2 13.2 14.2 37.0 38.2 0.7
-DoubleDipper 38.5 43.9 67.6 36.8 35.2 39.2 30.1 16.4
-Mistral-7b-instruct ICL 52.1 68.5 79 67.8 39.2 49.4 36.7 24.1
-DoubleDipper 52.8 58.9 95.6 82.8 44.4 43.1 26.4 18.4
-Table 5: Comparison of traditional In-Context Learning (ICL) where each demonstration example comprises a full text, a question and an answer to DoubleDipper where the demonstrations contain only question-answer pairs, automatically generated on the same input text.
-To analyze the performance gain of DoubleDipper compared to the traditional In-Context-Learning, we prompt a representative model from each family (Gemini Pro, Gemma 2B, Llama 7B, and Mistral) with few-shot examples from the same dataset, each example including its own context and question-answer pair.
-We report the results in Table 5.
-For all models, DoubleDipper achieves a higher performance than the traditional ICL.
-Interestingly, for Gemma 2B and Llama 7B, the traditional ICL not only underperforms DoubleDipper but it also significantly hurts performance compared to the baseline (e.g., -14.3 for Gemma 2B and -14.9 for Llama 7b).
-This performance drop might be because each demonstration in the traditional ICL setup comprises a few thousand tokens, which exacerbates the models’ existing challenges with processing long-range dependencies.
+6 Ablation Studies
+Gemma 2B Self Gemini Pro
+Gemini Pro 70.4 71.6 71.6
+Gemini Nano 62.1 61.7 62.8
+Gemini Flash 66.1 67.5 68.1
+Gemma 2B 49.5 49.5 51.0
+Gemma 9B 61.2 62.0 63.5
+Llama 3.1 59.9 60.1 61.5
+Mistral v0.3 51.0 49.8 52.6
+Mixtral 52.2 49.8 54.4
+Table 5: Average performance of DoubleDipper with different models for generating the demonstrations. See Appendix C.2 for the results on each evaluation dataset.
+How many examples are needed?
+In Table 6, we explore the impact of varying $k$, the number of self-generated few-shot examples in DoubleDipper to 1, 3, 5, and 10.
+On average, a single demonstration already provides an improvement over the baseline.
+Adding 3 demonstrations adds another boost of 2 points, while increasing the number of demonstrations to 5 and 10 leads to a marginal improvement.
+This finding is in line with previous work (Brown et al., 2020a; Min et al., 2022b).
+We conclude that a small number of examples carries most of the benefit with our method, but given additional computation budget, adding more examples does carry additional minor benefit.
+$k=1$ $k=3$ $k=5$ $k=10$
+Gemini Nano 60.0 62.1 62.2 62.3
+Gemini Flash 65.3 66.1 65.9 66.1
+Gemma 2B (v2) 47.0 49.5 49.6 49.9
+Gemma 9B (v2) 58.7 61.2 61.4 61.3
+Llama 3.1 57.7 59.9 60.6 61.4
+Mistral 7B (v0.3) 48.9 51.0 51.1 51.4
+Mixtral 7x8B (v0.1) 49.3 52.2 51.7 52.2
+Table 6: Average performance on our evaluation set with various numbers of self-generated few shot demonstrations ($k$) in DoubleDipper. See Appendix C.1 for the results on each evaluation dataset.
+Investigating the effect of the few-shot generator
+To understand the impact of the default chosen model (Gemma 2 2B) for generating the demonstrations, we conducted two additional experiments.
+The first experiment is Self in which we use the same model for generating the demonstrations and for answering the original question.
+In the second experiment, we generate the demonstrations with the best LLM used in our experiments, namely Gemini Pro.
+The average results are reported in Table 5 and the performance for each evaluation dataset is presented in Appendix C.2.
+The results show that generating the demonstrations with Gemma 2 or Self achieves similar performance, while Gemini Pro leads to a consistent increase in performance across models, indicating that future better models can improve further the performance.
 As mentioned in Section 3, DoubleDipper promotes also efficiency by adding to the original prompt only a few extra tokens, leading to a significantly cheaper inference than the traditional ICL.
-How many examples are needed?
-In Table 6, we explore the impact of varying $k$, the number of self-generated few-shot examples in DoubleDipper to 3, 5, and 10.
-Our analysis reveals no significant differences in performance across these variations, while 3 self-generated examples are sufficient to improve performance.
-This finding is in line with previous work (Brown et al., 2020a; Min et al., 2022b).
-$k=3$ $k=5$ $k=10$
-Gemini Pro 54.6 54.1 54.7
-Gemini Ultra 54.7 54.1 53.8
-Gemma-2b-it (v1.1) 41.2 40.8 41.0
-Gemma-7b-it (v1.1) 51.3 50.2 49.5
-Mistral-7b-instruct (v0.2) 52.8 52.8 52.4
-Llama-2-7b-chat 33.3 32.5 31.5
-Llama-2-13b-chat 34.9 34.4 33.0
-Llama-2-70b-chat 51.6 51.1 51.5
-Table 6: Average performance on our evaluation set with various numbers of self-generated few shot demonstrations ($k$) in DoubleDipper.
 DoubleDipper without identification of supporting paragraphs
-To ablate the second principle in DoubleDipper, namely the explicit identification of the supporting paragraphs before generating the answer, we prompt Gemini Pro and Gemini Ultra with self-generated few shot examples that comprise only question-answer pairs (without instructing the model to retrieve the relevant passage(s)).
-Gemini Pro achieves on average 36.2 points (-18.4) and Gemini Ultra achieves 54.0 (-0.7 points).
-This indicates that instructing small models (e.g., Gemini Pro) with explicit prediction of supporting paragraphs achieves similar results to large models (Gemini Ultra).
-Do few-shot examples really “instruct” the model to comprehend the text or provide clues to the query?
-To answer this question, we prompt Gemini Pro and Gemini Ultra with only the self-generated demonstrations without the input context.
-As expected, removing the input context leads to a huge performance drop, -31.7 for Gemini Pro and -22.8 for Gemini Ultra compared to DoubleDipper.
-This confirms that DoubleDipper indeed “teaches” the model how to comprehend the input.
-Qualitative Analysis
-We manually analyze 50 prompts, with few-shot demonstrations produced by Gemini Ultra.
+To ablate the second principle in DoubleDipper, which is the explicit identification of the supporting paragraphs before generating the answer, we prompt the open source models with self-generated few shot examples that comprise only question-answer pairs (without instructing the model to retrieve the relevant passage(s)).
+For all models, the average QA performance of DoubleDipper without the evidence is lower than DoubleDipper with evidence identification.
+When averaging results across models and datasets, we report a substantial F1 drop from 54.8 to 46.6.
+Full results are available in Appendix C.3 in Table 9.
+Qualitative analysis: Correctness of the generated QA pairs
+We manually analyze 150 QAs generated by Gemma 2B as demonstrations.
 Our review confirms that 93.5% of these self-generated QAs are correct, meaning that the question is meaningful and the answer could be found in the corresponding paragraph.
@@
 By recycling the input context to generate the few shot examples, each demonstration includes solely a question, an answer and a pointer to the relevant paragraph, without a separate context, thus effectively addressing the challenging of In-Context-Learning with long context.
-Experimental results show that our approach ourperforms both the vanilla LLM and the traditional ICL in various QA settings, including distractor passages in the input, True/False questions and multi-hop QA.
+Experimental results show that our approach substantially ourperforms the baselines in various QA settings, including distractor passages in the input, True/False questions and multi-hop QA.
 8 Limitations
@@
 basicstyle=,xleftmargin=2.8em,xrightmargin=1.5em, showstringspaces=false, showspaces=false, showtabs=false, tabsize=2, breaklines=true, flexiblecolumns=true, escapeinside=<@@>, breakatwhitespace=true
-Figure 5: Template prompt for the generation question-answer pairs.
-Appendix B Identification of Supporting Passages
-Table 7 presents the F1 results of the tested models for the supporting relevant passages identification.
-Without any demonstration (zero-shot), all open source models achieve a poor performance, while DoubleDipper significantly improves performance (+37.8 F1 for Llama2 70B, +38 F1 for Gemma 7B, +30.5 F1 for Mistral).
-Avg. Lost PIR MonoRel SRT HotPotQA 2Wiki MuSique
-Gemini Pro Baseline - - - - - - - -
-Zero-shot 54.5 58.2 47.1 75.9 46.6 56.3 58.3 39.4
-DoubleDipper (Self) 50.0 59.7 77.4 38.8 54.2 51.2 34.6 34.1
-DoubleDipper (PaLM 2) 50.6 60.1 47.3 68.0 21.2 58.8 57.9 41.0
-Gemini Ultra Baseline - - - - - - - -
-Zero-shot 40.4 57.4 30.3 39.4 17.0 52.4 43.7 42.9
-DoubleDipper (Self) 37.5 58.6 25.5 21.5 11.1 55.4 46.4 44.3
-DoubleDipper (PaLM 2) 38.4 58.0 28.0 24.6 12.0 55.3 47.0 43.7
-Gemma-2b-it (v1.1) Baseline - - - - - - - -
-Zero-shot 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
-DoubleDipper (Self) 23.3 10.4 43.1 37.5 13.5 20.1 13.0 25.3
-DoubleDipper (PaLM 2) 25.6 9.1 48.7 48.7 15 18.7 14 25.3
-Gemma-7b-it (v1.1) Baseline - - - - - - - -
-Zero-shot 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
-DoubleDipper (Self) 38.0 39.7 44.8 59.8 7.1 41.8 36.5 36.1
-DoubleDipper (PaLM 2) 36.9 39.2 45.2 58.6 7.3 40.1 32.8 35
-Mistral-7b-instruct (v0.2) Baseline - - - - - - - -
-Zero-shot 11.5 6.7 0.7 2.9 0.0 17.2 18.2 34.5
-DoubleDipper (Self) 42.0 48.2 66.2 60.6 5.8 41.7 29.3 42.2
-DoubleDipper (PaLM 2) 41.5 48.3 70.0 58.6 6.8 40.0 26.0 40.6
-Llama-2-7b-chat Baseline - - - - - - - -
-Zero-shot 0.8 0.2 0.0 0.0 3.2 0.0 0.0 2.1
-DoubleDipper (Self) 19.4 17.5 27.7 20.2 5.7 21.7 19.7 23.3
-DoubleDipper (PaLM 2) 19.6 15.8 29.3 21.6 5.7 22.2 20.7 22.0
-Llama-2-13b-chat Baseline - - - - - - - -
-Zero-shot 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
-DoubleDipper (Self) 28.8 16.5 51.0 32.0 18.4 29.6 22.2 31.7
-DoubleDipper (PaLM 2) 30.3 13.4 51.6 30.5 15.6 36.4 27.3 36.9
-Table 7: Performance (F1) of supporting paragraph(s) prediction.
-The baseline does not predict supporting paragraphs.
-DoubleDipper provides a significant performance boost for all open source models, whose performance is close to 0 (except from Mistral) in the zero shot setting.
+Figure 5: Template prompt for generating the QA pairs in DoubleDipper.
+Appendix B Lost-in-the-middle
+Figure 6 shows the QA accuracy of the models Gemma 2 2B, Mistral 7B, Gemini Nano and Gemini Pro on our subset of the “Lost-in-the-middle” dataset.
+Figure 6: Performance (accuracy) of Gemma 2 2B, Mistral 7B, Gemini Nano and Gemini Pro with and without DoubleDipper on our sample of the Lost-in-the-middle dataset (Liu et al., 2023) according to the position of the document that contains the answer.
+Appendix C Analysis
+C.1 Impact of the Number of Demonstrations in DoubleDipper
+Avg. 2Wiki MonoRel PIR SRT HotPotQA Lost MuSique
+Gemini Nano $k=1$ 60.03 37.68 85.20 88.20 56.40 62.50 60.68 29.56
+$k=3$ 62.12 40.55 86.60 95.40 56.20 65.12 60.52 30.44
+$k=5$ 62.25 41.79 87.00 95.60 55.20 65.05 60.56 30.52
+$k=10$ 62.33 43.16 86.00 96.20 55.20 65.37 60.44 29.96
+Gemini Flash $k=1$ 65.34 48.83 84.80 90.80 66.40 60.17 65.16 41.19
+$k=3$ 66.11 48.03 85.80 95.00 68.60 60.58 65.04 39.71
+$k=5$ 65.87 47.49 87.20 95.00 66.60 61.13 64.20 39.48
+$k=10$ 66.08 46.13 86.20 95.20 69.00 62.03 63.80 40.18
+Gemma 2B (v2) $k=1$ 47.05 24.05 73.60 77.00 50.20 41.32 47.04 16.13
+$k=3$ 49.48 23.66 85.80 81.60 50.00 39.85 46.68 18.77
+$k=5$ 49.59 26.70 85.40 80.40 48.40 40.34 46.56 19.30
+$k=10$ 49.92 26.43 85.80 81.40 49.00 40.91 47.68 18.26
+Gemma 9B (v2) $k=1$ 58.77 37.58 81.00 87.60 51.60 58.46 63.44 31.68
+$k=3$ 61.20 41.70 84.00 95.00 51.40 61.21 61.80 33.31
+$k=5$ 61.40 44.42 84.20 94.80 50.00 62.99 60.76 32.60
+$k=10$ 61.34 43.78 84.40 95.80 50.40 63.01 59.96 32.05
+Llama 3.1 8B $k=1$ 57.74 38.20 85.40 84.80 51.80 57.50 60.96 25.49
+$k=3$ 59.86 38.75 91.20 90.60 51.00 59.29 58.12 30.09
+$k=5$ 60.58 41.56 89.80 91.40 51.00 61.31 57.68 31.28
+$k=10$ 61.41 44.68 89.00 91.00 51.00 63.88 56.36 33.98
+Mistral 7B $k=1$ 48.91 28.86 67.60 77.40 50.00 40.62 62.00 15.91
+$k=3$ 51.03 28.65 68.40 88.80 50.60 43.44 60.68 16.66
+$k=5$ 51.12 29.86 69.20 89.00 49.40 43.83 60.36 16.23
+$k=10$ 51.44 30.14 68.20 89.40 49.00 46.49 59.96 16.91
+Mixtral 7x8B $k=1$ 49.26 19.83 86.00 74.80 49.40 31.73 69.00 14.07
+$k=3$ 52.22 22.31 91.80 86.00 47.80 35.10 66.56 15.97
+$k=5$ 51.65 23.55 91.40 82.40 47.20 34.64 64.92 17.46
+$k=10$ 52.18 23.49 91.80 84.40 47.00 36.23 64.00 18.32
+Table 7: Performance of DoubleDipper on our evaluation set with various numbers of self-generated few shot demonstrations ($k$).
+Table 7 presents the results of DoubleDipper with 1, 3 (main experiment in the paper), 5 and 10 generated demonstrations. For all these experiments, the demonstrations were generated by Gemma 2 2B.
+Figure 7 shows the QA accuracy of DoubleDipper on “Lost” according to the position of the relevant passage for each $k\in\{1,3,5,10\}$.
+Figure 7: Performance (accuracy) of several models with DoubleDipper according to the number of self-generated demonstrations in the prompt.
+C.2 Impact of the few-shot generator
+Table 8 presents the detailed QA performance of all models with different models for generating DoubleDipper’s demonstrations.
+As mentioned in the paper (Section 6), generating the demonstrations with the best model (ie.
+Gemini Pro) achieves the best performance overall.
+Avg. 2Wiki MonoRel PIR SRT HotPotQA Lost MuSique
+Gemini Pro Gemma 2B 70.4 46.8 97.4 99.0 79.6 60.9 72.4 36.4
+Self 71.6 49.6 98.2 99.6 78.8 62.5 72.5 40.1
+Gemini Flash Gemma 2B 66.1 48.0 85.8 95.0 68.6 60.6 65.0 39.7
+Self 67.5 49.5 88.6 96.8 65.6 64.3 65.2 42.5
+Gemini Pro 68.1 50.3 90.4 96.0 69.0 64.8 64.8 41.6
+Gemini Nano Gemma 2B 62.1 40.6 86.6 95.4 56.2 65.1 60.5 30.4
+Self 61.7 39.8 85.8 95.2 56.2 64.3 60.7 30.0
+Gemini Pro 62.8 40.4 87.2 99.0 54.4 66.8 60.4 31.7
+Gemma 2 2B Self 49.5 23.7 85.8 81.6 50.0 39.9 46.7 18.8
+Gemini Pro 51.0 25.2 83.8 93.4 48.0 41.7 46.2 18.4
+Gemma 2 9B Gemma 2B 61.2 41.7 84.0 95.0 51.4 61.2 61.8 33.3
+Self 62.0 44.8 82.4 97.0 50.6 62.9 61.5 34.9
+Gemini Pro 63.5 46.2 86.4 99.2 52.2 64.4 61.4 34.8
+Llama 3.1 8B Gemma 2B 59.9 38.7 91.2 90.6 51.0 59.3 58.1 30.1
+Self 60.1 41.3 88.6 89.8 50.8 61.5 57.8 31.0
+Gemini Pro 61.5 41.0 89.8 95.6 51.6 63.5 57.5 31.7
+Mistral 7B (v0.3) Gemma 2B 51.0 28.6 68.4 88.8 50.6 43.4 60.7 16.7
+Self 49.8 25.2 68.8 93.8 49.6 36.5 60.6 14.3
+Gemini Pro 52.6 31.4 67.2 96.4 48.2 45.7 62.1 17.0
+Mixtral 7x8B (v0.1) Gemma 2B 52.2 22.3 91.8 86.0 47.8 35.1 66.6 16.0
+Self 49.8 16.0 91.8 86.0 48.8 25.5 67.8 12.8
+Gemini Pro 54.4 20.8 94.8 96.2 49.0 36.3 66.6 16.9
+Table 8: Performance of DoubleDipper according to the model for generating the demonstrations (Gemma 2B, Self or Gemini Pro).
+C.3 Impact of the identification of supporting paragraphs in the QA generation
+Table 9 compares the performance of DoubleDipper to DoubleDipper without evidence identification.
+Avg. 2Wiki MonoRel PIR SRT HotPotQA Lost MuSique
+Gemma 2B DoubleDipper (QA only) 46.4 16.0 84.6 79.2 49.6 31.1 49.5 14.7
+DoubleDipper 49.5 23.7 85.8 81.6 50.0 39.9 46.7 18.8
+Gemma 9B DoubleDipper (QA only) 55.1 22.5 91.0 96.6 52.0 42.9 61.0 20.0
+DoubleDipper 61.2 41.7 84.0 95.0 51.4 61.2 61.8 33.3
+Mistral 7B (v0.3) DoubleDipper (QA only) 49.1 21.8 73.4 89.8 48.6 38.6 60.3 11.5
+DoubleDipper 51.0 28.6 68.4 88.8 50.6 43.4 60.7 16.7
+Mixtral 8x7B (v0.1) DoubleDipper (QA only) 49.5 17.0 91.4 86.8 50.8 22.9 65.8 11.9
+DoubleDipper 52.2 22.3 91.8 86.0 47.8 35.1 66.6 16.0
+Llama 3.1 8B DoubleDipper (QA only) 32.7 25.0 9.2 41.4 46.6 33.5 56.7 16.6
+DoubleDipper 59.9 38.7 91.2 90.6 51.0 59.3 58.1 30.1
+Table 9: Performance of DoubleDipper and DoubleDipper without instructing the models to retrieve the evidence (QA only) on the QA datasets.

### v4

@@
-Can Few-shot Work in Long-Context? Recycling the Context to Generate Demonstrations
+DoubleDipper: Improving Long-Context LLMs via Context Recycling
 Abstract
 Despite recent advancements in Large Language Models (LLMs), their performance on tasks involving long contexts remains sub-optimal.
-In-Context Learning (ICL) with few-shot examples may be an appealing solution to enhance LLM performance in this scenario; However, naïvely adding ICL examples with long context introduces challenges, including substantial token overhead added for each few-shot example and context mismatch between the demonstrations and the target query.
-In this work, we propose to automatically generate few-shot examples for long context QA tasks by recycling contexts.
+In this work, we propose DoubleDipper, a novel In-Context-Learning method that automatically generates few-shot examples for long context QA tasks by recycling contexts.
 Specifically, given a long input context (1-3k tokens) and a query, we generate additional query-output pairs from the given context as few-shot examples, while introducing the context only once.
@@
 We further enhance each demonstration by instructing the model to explicitly identify the relevant paragraphs before the answer, which improves performance while providing fine-grained attribution to the answer source.
-We apply our method on multiple LLMs and obtain substantial improvements (+16 absolute points on average across models) on various QA datasets with long context, especially when the answer lies within the middle of the context.
-Surprisingly, despite introducing only single-hop ICL examples, LLMs also successfully generalize to multi-hop long-context QA using our approach.
+We apply our method on multiple LLMs and obtain substantial improvements (+16 absolute points on average across models) on various QA datasets with long context.
+Surprisingly, despite introducing only single-hop ICL examples, LLMs successfully generalize to multi-hop long-context QA using our approach.
 […]
 tcb@breakable
-Can Few-shot Work in Long-Context?
-Recycling the Context to Generate Demonstrations
+DoubleDipper: Improving Long-Context LLMs via Context Recycling
 Arie Cattan1,2thanks: Work done during an internship at Google.
@@
 1 Introduction
-Figure 1: Performance of Gemini Flash (v1.5) on a sample of the Lost-in-the-middle dataset (Liu et al., 2023).
-The X-axis is the position of the relevant passage in the context.
-The baseline (blue line) displays a U-shaped curve, performing well only when the relevant passage is at the beginning or end of the input.
-The oracle (green line) shows significant performance gain when the relevant passage ID is provided in the prompt, showing that the identification of supporting evidence(s) is a major challenge.
-DoubleDipper (our method, orange line) flattens this U-shaped trend.
 Long contexts are prevalent in various domains, ranging from legal documents and scientific articles to lengthy reports and novels.
-These may consist of a single extensive document or multiple passages, typically retrieved through specific retrieval mechanisms (e.g., RAG; Lewis et al., 2020).
+These may consist of a single extensive document or multiple passages, typically retrieved through specific retrieval mechanisms (e.g., RAG Lewis et al.
+(2020)).
+Figure 1: Comparison of traditional In-Context-Learning (ICL) and our new method.
+In traditional ICL (left), each example comprises a possibly lengthy context, accompanied by a query and an answer, typically derived from the training dataset.
+Conversely, our approach (right) simplifies each example to just a question and an answer, both of which are generated directly from the provided input context.
 Yet, while Large Language Models (LLMs) have demonstrated impressive capabilities in a variety of tasks including answering questions requiring one or multiple reasoning steps, they often struggle to answer simple questions when faced with long contexts.
-Despite substantial engineering efforts (Chen et al., 2023) to extend the context window of LLMs to extremely long inputs (32k and even 1M tokens), these models continue to struggle with much shorter inputs, comprising only a few thousand tokens.
-In order to answer questions from long inputs, models should implicitly identify relevant information segments and then reason over these segments to formulate an answer.
-It has been shown that LLMs struggle when the relevant information is buried in the middle of the context (Liu et al., 2023) or obscured by numerous irrelevant details (Levy et al., 2024).
-Our analysis (illustrated in Figure 1) identifies the identification of relevant information as a major performance bottleneck of current models in long contexts.
-In this work, we introduce a novel method to enhance the QA performance of LLMs in long input setups (to allow direct comparisons across a wide swath of models, we limit "long context" here to 1-3k tokens).
-Our approach, termed DoubleDipper, leverages LLMs’ In-Context Learning (ICL) capability and is based on two principles.
-First, instead of typical ICL, where each few-shot example is standalone with a separate length context and a question-answer (QA) pair, we propose to recycle the given input context and automatically generate few-shot examples from this context.
+Despite substantial engineering efforts (Chen et al., 2023b) to extend the context window of LLMs to extremely long inputs (32k and even 1M tokens), these models continue to struggle with much shorter inputs, comprising only a few thousand tokens, especially when the relevant information is buried in the middle of the context (Liu et al., 2023b) or obscured by numerous irrelevant details (Levy et al., 2024).
+In-Context Learning (ICL) with few-shot examples may be an appealing solution to enhance LLM performance in long contexts.
+However, applying ICL in real-world scenarios without access to training data introduces significant challenges.
+Developers need to maintain a demonstration pool for retrieving semantically similar demonstrations to any given query (Liu et al., 2022; Rubin et al., 2022).
+Furthermore, ICL adds a substantial token overhead to the prompt, an issue that becomes even more pronounced with long-context inputs.
+In this work, we introduce a novel method to enhance the QA performance of LLMs in long input setups.
+Our approach, termed DoubleDipper, leverages LLMs’ In-Context Learning capability and is based on two principles.
+First, instead of typical ICL, where each few-shot example is standalone with a separate lengthy context and a question-answer (QA) pair, we propose to recycle the given input context and automatically generate few-shot examples from this context.
 Specifically, we randomly select a few paragraphs from the given input context and generate QA pairs for each passage.
@@
 Figure 1 illustrates the differences between the traditional ICL with few-shot examples and DoubleDipper.
-Second, we further enrich each ICL example to instruct the model to explicitly identify the paragraph(s) containing the relevant information before generating the answer.
-This can be regarded as a structured Chain of Thought that incentivizes the model to pinpoint relevant information before reasoning, an essential capability for long-context processing.
-Figure 2: Comparison of traditional In-Context-Learning (ICL) and our new method.
-In traditional ICL (left), each example comprises a possibly lengthy context, accompanied by a query and an answer, typically derived from the training dataset.
-Conversely, our approach (right) simplifies each example to just a question and an answer, both of which are generated directly from the provided input context.
+Second, we enhance each ICL demonstration by explicitly instructing the model to identify relevant information prior to generating an answer.
+Explicitly identifying relevant passages can be regarded as a structured Chain of Thought that incentivizes the model to pinpoint relevant information before reasoning, an essential capability for long-context processing.
 By generating few-shot demonstrations from various sections of the input context while instructing the model to identify relevant passages, DoubleDipper encourages the model to develop deeper reading comprehension skills specific to the given input evidence.
@@
 DoubleDipper presents several advantages.
-In terms of efficiency, since each example does not include its own input context, our method adds to the original prompt a minimal number of tokens, resulting in a substantially cheaper inference than traditional ICL.
-Additionally, recycling the same context for ICL demonstrations ensures that the few-shot examples refer to the same domain as the input question, thus obviating the need for external retrieval processes.
-Finally, DoubleDipper generates answers with attribution to relevant paragraphs, improving the model’s lookup ability and offering transparency, which substantially simplifies human evaluation (Slobodkin et al., 2024).
-We applied DoubleDipper to a variety of LLMs, both commercial (Gemini Pro, Nano and Flash; Reid et al., 2024) and open-source (Llama 3.1 (Dubey et al., 2024), Mistral (Jiang et al., 2023), Mixtral (Jiang et al., 2024) and Gemma (Riviere et al., 2024)), and evaluate it on various QA datasets with long inputs, including common multi-hop QA datasets.
+First, recycling the same context for ICL demonstrations ensures that the few-shot examples refer to the exact same domain as the input question, thus obviating the need for external retrieval of similar demonstrations.
+Also, in terms of efficiency, since each example does not include its own input context, our method adds to the original prompt a minimal number of tokens, resulting in a substantially cheaper inference than traditional ICL.
+Finally, DoubleDipper generates answers with attribution to relevant paragraphs, improving the model’s lookup ability and offering transparency, which substantially simplifies human evaluation (Menick et al., 2022; Gao et al., 2023; Liu et al., 2023a; Slobodkin et al., 2024).
+We applied DoubleDipper to 12 LLMs, both commercial (Gemini Pro, Nano, Flash (Reid et al., 2024); GPT-4 (Achiam et al., 2023)) and open-source ranging from 2B to 70B parameters (Llama 3.1 (Dubey et al., 2024); Mistral (Jiang et al., 2023); Mixtral (Jiang et al., 2024); Gemma, (Riviere et al., 2024)).
+We evaluate our method on 7 QA datasets with long inputs, including common multi-hop QA datasets.
 Our experiments demonstrate that with only 3 self-generated few-shot examples, DoubleDipper consistently outperforms the baseline on our evaluation set by 16 absolute points on average across models.
@@
 Challenges in Long Context for Language Modeling.
-LLMs have been well-documented to struggle when input length grows (An et al., 2023), and especially so when it exceeds input lengths seen during training (Anil et al., 2022).
-Various methods have been proposed to advance long-context capabilities: Architectural, e.g., to augment the embedding layer to cleverly extrapolate to unseen lengths (Vaswani et al., 2017; Press et al., 2021; Caciularu et al., 2022; Hsieh et al., 2024); and via data, e.g., to incorporate longer inputs and more challenging long-context scenarios into training (He et al., 2023; Chen et al., 2023).
-However, this challenging problem stubbornly remains in competitive models today (Liu et al., 2023; Bishop et al., 2023; Levy et al., 2024).
-In contrast to the above methods, DoubleDipper does not involve training or architectural changes.
-In documenting and exploring LLM performance in long-context settings, many different benchmarks targeting it have been proposed, such as Scrolls and Zero-Scrolls (Shaham et al., 2022, 2023), Loogle (Li et al., 2023), LongBench (Bai et al., 2023), inter alia.
+LLMs have been well-documented to struggle when input length grows (An et al., 2023), and especially when it exceeds input lengths seen during training (Anil et al., 2022).
+Various methods have been proposed to advance long-context capabilities: Architectural, e.g., to augment the embedding layer to cleverly extrapolate to unseen lengths (Vaswani et al., 2017; Press et al., 2021; Caciularu et al., 2022; Tan et al., 2024); via data, e.g., to incorporate longer inputs and more challenging long-context scenarios into training (Chen et al., 2024b; He et al., 2024; Chen et al., 2024a); via attention intervention (Hsieh et al., 2024) or by considering question likelihood as a signal for prompt reordering (Liu et al., 2024b).
+However, this challenging problem stubbornly remains in competitive models (Liu et al., 2023b; Bishop et al., 2024; Levy et al., 2024).
+In contrast to the above methods, DoubleDipper is a simple method that does not involve training or architectural changes.
+Many benchmarks targetting long-context have been proposed, such as Scrolls and Zero-Scrolls (Shaham et al., 2022, 2023), Loogle (Li et al., 2023), LongBench (Bai et al., 2023b), L-Eval (An et al., 2024), inter alia.
 The problem of designing informative and reliable benchmarks in long-context is an an active, ever-changing area of research (Goldman et al., 2024; Yen et al., 2024).
@@
 In-Context Learning
-The area of in-context learning (ICL) is a class of prompting techniques where demonstrations are added to the prompt in order to steer or improve model behavior (Min et al., 2022a).
-Typically, in-context learning involves hand-crafted demonstrations (Song et al., 2022), automatic retrieval of demonstrations from a larger set (Paranjape et al., 2023), or instructing the model to perform various tasks one after another in a pipeline (Gao et al., 2022).
-Recent improvements in long-context capabilities of LLMs have also had effect on improving the yield from in-context learning by simply using more short-length demonstrations (Agarwal et al., 2024).
-While such methods are widely used for their effectiveness (Brown et al., 2020b; an Luo et al., 2024), they remain under-explored in settings of long-context.
-The reason is simple: If the context is already extremely long, adding additional demonstrations comparable in length to the input context will likely amplify the existing limitations of long context handling (Li et al., 2024c).
-More related to our work, a few recent studies propose to prompt LLMs for automatically generating in-context demonstrations for various short context tasks (Kim et al., 2022; Yasunaga et al., 2024; Li et al., 2024a, b).
-However, applying those methods to long contexts is not straightforward because LLMs would need to generate long demonstrations.
-In this work, we tackle these challenges and present a novel ICL method that recycles the given long context and further instructs models to identify the relevant information before generating the answer.
+In-context Learning (ICL) consists of adding demonstrations to the prompt in order to steer or improve model behavior (Min et al., 2022).
+These demonstrations are either hand-crafted (Song et al., 2022), or retrieved from a large set of training examples (Liu et al., 2022; Rubin et al., 2022; Paranjape et al., 2023).
+While ICL provides a flexible approach to learn new tasks without updating parameters (Brown et al., 2020b; an Luo et al., 2024), applying ICL in real-world scenarios is challenging notably because there is no available training data for each user query.
+More closely related to our work, a few recent studies propose to prompt LLMs for automatically generating in-context demonstrations for various short context tasks.
+For instance, Kim et al., 2022 focus on sentence classification tasks and prompt LLMs to generate full demonstrations conditioned on a label (e.g., “write a negative review”) and (Chen et al., 2023a; Yasunaga et al., 2024; Li et al., 2024) generate relevant exemplars to the query for reasoning problems.
+While effective for reasoning tasks with short contexts, these methods are not directly applicable to long-context scenarios because LLMs would need to generate not just a question and an answer, but also an entire long context for each demonstration, which is both computationally expensive and prone to hallucination.
+DoubleDipper addresses these challenges by introducing a novel and efficient strategy: rather than generating entirely new contexts, it recycles the input context and generates only the question-answer (QA) pairs needed as demonstrations.
 3 DoubleDipper
@@
 As each QA is associated with its respective paragraph, we form the demonstrations to instruct the model to identify the relevant passage(s) and the correct answer.
-Recall that our work focuses on the task of question answering (QA) with long input context comprising multiple paragraphs.
-In addition to the answer, we aim to identify the supporting paragraphs in order to provide attribution.
-Formally, given a long input text $D$ composed of $n$ paragraphs $D=\{p_{1},p_{2},...,p_{n}\}$ and a question $q$, the goal is to generate the answer $a$ and identify the set(s) of paragraphs that support the answer $S=\{s_{1},...,s_{k}\}$.
+Given a long input text $\mathcal{C}$ composed of $n$ paragraphs $\mathcal{C}=\{p_{1},p_{2},...,p_{n}\}$ and a question $q$, the goal is to generate the answer $a$ and identify the set(s) of paragraphs that support the answer $S=\{s_{1},...,s_{k}\}$.
 The number of the supporting paragraphs is not known in advance and can be one or more.
@@
 Figure 2 illustrates DoubleDipper.
-Starting with the input paragraphs $D$, we initially select $k$ paragraphs at random (e.g., paragraphs $15$, $5$, and $17$, for $k:=3$).
-For each chosen paragraph, we prompt the model to formulate a question that pertains to the specific paragraph, accompanied by an appropriate answer (for further details on prompt specifications, refer to Appendix A).
+Starting with the input paragraphs $\mathcal{C}$, we initially select $k$ paragraphs at random (e.g., paragraphs 151515, 555, and 171717, for $k:=3$).
+For each chosen paragraph, we prompt the model to formulate a question that pertains to the specific paragraph, accompanied by an appropriate answer (for further details on prompt specifications, refer to Appendix B).
 Each generated QA pair is directly associated with its origin paragraph, enabling us to assemble the following structured in-context demonstration, shown as the DoubleDipper block in Figure 2:
@@
 Here, $p_{i}$ indicates the index of the paragraph associated with the QA pair $(q_{i},a_{i})$.
-Given a test question $q$, we then form a QA prompt by concatenating the original input context $D$, the compiled demonstrations and $q$.
-The model first generates the one or more indices of the supporting paragraph(s), followed by the answer.
-Unlike traditional few shot examples that instruct the model about a specific task, DoubleDipper aims to coach the model on how to “handle” the specific input context.
+Given a context $\mathcal{C}$ and a test question $q$, we compile a list of generated demonstrations $\mathcal{D}_{\text{demo}}=(q_{1},p_{1},a_{1},\dots,q_{k},p_{k},a_{k})$ to predict the output $y\sim p_{\theta}(y\mid\mathcal{C},\mathcal{D}_{\text{demo }},q)$ where the output $y$ is the concatenation of one or more indices of the supporting paragraph(s) $S$ and the answer $a$.
+Unlike traditional few shot examples that instruct the model about a specific task, DoubleDipper aims to coach the model on how to “handle” the input context.
 This is achieved by guiding the model to explicitly localize relevant information before generating the answer.
 Also, by randomly sampling multiple paragraphs from the input, DoubleDipper guarantees that the ICL demonstrations involve reading different parts of the context, allowing the model to better comprehend the input text.
-Beyond improving the performance of the QA task, instructing the model to provide the supporting paragraphs is valuable on its own as it offers transparency and substantially eases human evaluation (Slobodkin et al., 2024).
+Beyond improving the performance of the QA task, instructing the model to provide the supporting paragraphs offers transparency and substantially eases human evaluation.
 DoubleDipper offers several advantages.
-First, as each example in the demonstration consists only of a question, an answer and the ID of relevant passage, the number of added tokens due to the extra demonstrations is minimal, leading to a low additional cost and computation compared to the traditional In-Context-Learning.
+First, as each example in the demonstration consists only of a question, an answer and the ID of relevant passage, the number of added tokens due to the extra demonstrations is minimal (5%), leading to a low additional cost and computation compared to the traditional In-Context-Learning.
 Furthermore, by reusing the same context to generate demonstrations, our approach guarantees that all few shot examples are derived from the exact same domain as the input query (Rubin et al., 2022).
-Finally, we observe that, although the QA pairs in the demonstration are confined to individual paragraphs, the actual query $q$ may require reasoning over multiple paragraphs (i.e., multi-hop QA).
-Surprisingly, LLMs can generalize from DoubleDipper local examples to these complex, global questions and successfully generate indices to multiple paragraphs (see Section 5).
 4 Experiments
-Size Context Window
-Gemini v1.5 Nano 32k
-Flash 32k
-Pro 32k
-Llama 3.1 7B 128k
-Mistral v0.3 7B 32k
-Mixtral v0.1 7x8B 32k
-Gemma 2 2B 8k
-9B 8k
-Table 1: Models used with DoubleDipper.
 Datasets
@@
 MuSiQue 500 2,549
-Table 2: Evaluation datasets in our experiments. The average number of tokens is computed according to Gemma’s tokenization of the simple prompt.
+Table 1: Evaluation datasets in our experiments. The average number of tokens is computed according to Gemma’s tokenization of the vanilla prompt.
 We apply our method to various datasets, each presenting its own domain-specific challenges.
@@
 Overall our evaluation set includes 5.5K instances, with statistics of each dataset given in Table 1.
-The Lost-in-the-middle dataset (Liu et al., 2023) includes examples from NaturalQuestions-Open (Kwiatkowski et al., 2019; Lee et al., 2019).
+The Lost-in-the-middle dataset (Liu et al., 2023b) includes examples from NaturalQuestions-Open (Kwiatkowski et al., 2019).
 Each instance consists of twenty Wikipedia passages, with only one passage containing the answer to the query.
 The remaining passages are distractors that are lexically similar but do not contain the answer.
-To assess the robustness of large language models (LLMs) to the position of relevant information, Liu et al.
-(2023) evaluated cases where the relevant passage appeared in positions 1, 5, 10, 15, and 20.
+To assess the robustness of LLMs to the position of relevant information, Liu et al.
+(2023b) evaluated cases where the relevant passage appeared in positions 1, 5, 10, 15, and 20.
 Following their methodology, we sampled 500 instances for each position, resulting in a total of 2,500 instances.
@@
 FLenQA includes three subtasks.
-The first subtask is Monotone Relations (MonoRel), where each instance asks whether a transitive relation between two entities holds based on the context (e.g., "Is X younger than Y?" based on the sentences "X is younger than Z" and "Z is younger than Y").
-The second subtask, People In Rooms (PIR), involves one key sentence indicating that a person is in a specific room and another key sentence describing a property of this room.
+The first subtask is MonoRel, where each instance asks whether a transitive relation between two entities holds based on the context (e.g., "Is X younger than Y?" based on the sentences "X is younger than Z" and "Z is younger than Y").
+The second subtask, PIR, involves one key sentence indicating that a person is in a specific room and another key sentence describing a property of this room.
 The question asks whether the person is in a room with the described property.
-The final subtask is Simplified Rule Taker (SRT), based on RuleTaker (Clark et al., 2020).
+The final subtask is SRT, based on RuleTaker (Clark et al., 2020).
 Each instance consists of a logical rule, two sentences each introducing a fact, and a question over the rule and facts.
 For each subtask, FLenQA includes contexts with varying lengths, from 50 to 3,000 tokens, by simply adding irrelevant text, demonstrating consistent performance degradation with increased input length.
-In our experiments, we sampled 250 instances for each subtask with input lengths of 2,000 and 3,000 tokens, leading to a total of 1,500 instances for FLenQA.
+In our experiments, we sampled 250 instances for each subtask with input lengths of 2,000 and 3,000 tokens, leading to a total of 1,500 instances.
 In addition, we evaluate our method on common multi-hop QA benchmarks.
-We sampled 500 instances from each of the following datasets: HotPotQA (Yang et al., 2018), 2Wiki (Ho et al., 2020), and MuSiQue (Trivedi et al., 2021).
+We sampled 500 instances from HotPotQA (Yang et al., 2018), 2Wiki (Ho et al., 2020), and MuSiQue (Trivedi et al., 2021).
 In all these datasets, the input text includes multiple passages, and models need to perform at least two steps of reasoning over different passages in order to answer the question.
+Models
+We apply DoubleDipper to a variety of models, both commercial and open-source.
+The commercial models include Gemini 1.5 Pro, Gemini 1.5 Flash (Reid et al., 2024) and GPT-4o-mini (Achiam et al., 2023).
+The open-source models we tested are Llama 3.1 8B, Llama 3.1 70B (Dubey et al., 2024), Gemma 2B (v2), Gemma 9B (v2) and Gemma 27B (Riviere et al., 2024), Mistral-7B-Instruct (v0.2) (Jiang et al., 2023), Mixtral-8x7B-Instruct (v0.1) (Jiang et al., 2024) and Mistral Nemo Instruct 2407111https://mistral.ai/news/mistral-nemo/.
+Few-shot generation in DoubleDipper is an auxiliary task and should ideally run in an efficient time without requiring heavy resources.
+Therefore, in our main experiments, we employ Gemma 2B to generate the demonstrations at it is the smallest and most efficient model used in our experiments.
+See Section 6 for an ablation analysis of the effect of the chosen model for generating the demonstrations.
+Baselines
+We evaluate DoubleDipper against two main baselines.
+The first is a vanilla baseline that takes as input the entire context $\mathcal{C}$ and the query $q$ and generates only the answer $a$, a common prompting strategy in recent studies on long context (Liu et al., 2023b; Levy et al., 2024).
+The second baseline, Zero-shot + Evidence Retrieval, prompts the model in a zero-shot manner to first identify relevant passages before generating the answer, following common practices in generating with attribution (Gao et al., 2023; Slobodkin et al., 2024; Fierro et al., 2024).
+Evaluation
+We evaluate each dataset with the original evaluation metrics.
+Namely, we report Accuracy for Lost-in-the-middle (Liu et al., 2023b) and FLenQA (Levy et al., 2024), and Token F1 for HotPotQA (Yang et al., 2018), 2Wiki (Ho et al., 2020) and MuSique (Trivedi et al., 2021).
+In addition to the task’s accuracy, we also evaluate the performance of the identification of the supporting paragraph(s), by computing the F1 score on the predicted set of supporting passages compared to the ground truth (Yang et al., 2018; Ho et al., 2020; Trivedi et al., 2021).
+Implementation Details
+We randomly select three passages from the input, each containing at least two sentences, and ask the model to generate a single QA pair for each passage.
+See Section 6 for an analysis of the number of self-generated demonstrations on the performance
+5 Results
 Avg. 2Wiki MonoRel PIR SRT HotPotQA Lost MuSique
@@
    DoubleDipper 66.1 48.0 85.8 95.0 68.6 60.6 65.0 39.7
-Gemini Nano (vanilla) 41.6 10.8 72.2 66.8 55.4 21.3 59.6 5.2
-   Zero-shot + Evidence Retrieval 56.5 32.0 82.4 82.4 56.4 56.7 60.5 25.2
-   DoubleDipper 62.1 40.6 86.6 95.4 56.2 65.1 60.5 30.4
-Gemma 2 2B (v2) (vanilla) 38.6 8.9 71.8 68.6 51.2 13.3 49.6 6.5
-   Zero-shot + Evidence Retrieval 42.0 22.3 66.8 70.6 40.2 30.6 47.6 16.2
-   DoubleDipper 49.5 23.7 85.8 81.6 50.0 39.9 46.7 18.8
 Gemma 2 9B (v2) (vanilla) 44.0 11.4 74.8 81.8 55.6 13.8 61.0 9.3
@@
    DoubleDipper 61.2 41.7 84.0 95.0 51.4 61.2 61.8 33.3
+Gemma 2 27B (v2) (vanilla) 48.0 11.2 85.2 79.2 63.8 14.3 62.4 19.9
+   Zero-shot + Evidence Retrieval 59.7 34.3 90.6 87.4 59.0 51.7 65.2 29.8
+   DoubleDipper 64.2 42.0 92.0 96.4 58.6 64.2 62.5 33.9
 Llama 3.1 8B (vanilla) 37.2 11.8 56.2 52.2 48.6 20.7 63.5 7.4
@@
    DoubleDipper 59.9 38.7 91.2 90.6 51.0 59.3 58.1 30.1
+Llama 3.1 70B (vanilla) 67.5 46.5 93.2 95.6 83.2 57.0 69.6 27.6
+   Zero-shot + Evidence Retrieval 71.0 57.6 97.8 97.6 81.0 62.0 71.5 29.5
+   DoubleDipper 72.9 62.3 98.6 98.8 73.0 71.5 66.0 40.2
 Mistral 7B (v0.3) (vanilla) 37.4 14.1 59.2 57.4 50.2 15.8 60.8 4.6
@@
    DoubleDipper 51.0 28.6 68.4 88.8 50.6 43.4 60.7 16.7
-Mixtral 7x8B (v0.1) (vanilla) 42.6 13.7 73.0 66.2 51.0 18.2 67.7 8.4
-   Zero-shot + Evidence Retrieval 47.4 18.8 81.8 73.6 50.6 26.3 67.9 13.1
-   DoubleDipper 52.2 22.3 91.8 86.0 47.8 35.1 66.6 16.0
+Mistral-Nemo 44.0 17.2 72.6 67.0 51.0 28.9 59.7 11.9
+   Zero-shot + Evidence Retrieval 46.7 29.1 59.8 67.2 51.0 39.8 60.6 19.4
+   DoubleDipper 53.3 38.7 58.0 81.0 51.4 51.9 62.9 29.5
 Table 2: Accuracy of the QA task for the vanilla baseline (prompting the model to only answer the question), Zero-shot + Evidence Retrieval (prompting the model to explicitly identify the relevant passage(s) before generating the answer) and DoubleDipper with 3 demonstrations generated by Gemma 2 2B.
-Models
-We apply DoubleDipper to a variety of models, both commercial and open-source.
-The commercial models include Gemini 1.5 Pro, Gemini 1.5 Flash and Gemini 1.0 Nano (Reid et al., 2024).
-The open-source models we tested are Llama 3.1 8B (Dubey et al., 2024), Gemma 2B (v2) and Gemma 9B (v2) (Riviere et al., 2024), Mistral-7B-Instruct (v0.2) (Jiang et al., 2023) and Mixtral-8x7B-Instruct (v0.1) (Jiang et al., 2024).
-Details about models’ size and context window are shown in Table 1.
-Few-shot generation in DoubleDipper is an auxiliary task and should ideally run in an efficient time without requiring heavy resources.
-Therefore, in our main experiments, we employ Gemma 2B to generate the demonstrations at it is the smallest and most efficient model used in our experiments.
-See Section 6 for an ablation analysis of the effect of the chosen model for generating the demonstrations.
-Baselines
-We compare DoubleDipper to the vanilla baseline, which takes as input the entire context $D$ and the query $q$ and generates only the answer $a$.
-We also compare to Zero-shot + Evidence Retrieval, which prompts the model in a zero-shot setting to identify the relevant passage(s) before generating the answer.
-Evaluation
-We evaluate each dataset with the original evaluation metrics.
-Namely, we report Accuracy for Lost-in-the-middle (Liu et al., 2023) and FLenQA (Levy et al., 2024), and Token F1 for HotPotQA (Yang et al., 2018), 2Wiki (Ho et al., 2020) and MuSique (Trivedi et al., 2021).
-In addition to the task’s accuracy, we also evaluate the performance of the identification of the supporting paragraph(s), by computing the F1 score on the predicted set of supporting passages compared to the ground truth (Yang et al., 2018; Ho et al., 2020; Trivedi et al., 2021).
-Implementation Details
-We randomly select three passages from the input (see Section 6 for an analysis of the number of self-generated demonstrations on the performance), each containing at least two sentences, and ask the model to generate a single QA pair for each passage (see Appendix A for the exact prompt).
-For all experiments, including few-shot generation and question-answering, we use a temperature setting of 0.
-5 Results
 Result 1: DoubleDipper offers a substantial performance boost.
 Table 2 presents the QA performance of the baseline, Zero-shot + Evidence Retrieval and DoubleDipper on our evaluation set.
+For brevity, we report the results of eight models here and four models in Appendix C in Table 7, which show similar trends.
 The results first show that prompting models to explicitly identify the relevant paragraphs before generating the answer (Zero-shot + Evidence Retrieval) leads to a performance improvement of 9.7 points on average across models over the vanilla baseline.
-DoubleDipper offers an additional substantial boost of 6.3 points for all models on average, culminating in a overall improvement of 16 absolute points over the vanilla baseline.
+DoubleDipper, leveraging demonstrations generated with the efficient Gemma 2B parameters, offers an additional substantial boost of 6.3 points for all models on average, culminating in an overall improvement of 16 absolute points over the vanilla baseline.
 Notably, while DoubleDipper produces simple QAs answerable from a single paragraph, it always surpasses the baseline in multi-hop QA datasets (HotPotQA, 2Wiki and MuSique).
-Likewise, DoubleDipper outperforms the baseline also on the FLenQA datasets (PIR, MonoRel and SRT), which involve synthetic True/False questions although the demonstrations in DoubleDipper are typically simple factoid questions.
+Likewise, DoubleDipper outperforms the baseline also on most FLenQA datasets (PIR, MonoRel and SRT), which involve synthetic True/False questions although the demonstrations in DoubleDipper are typically simple factoid questions.
+While DoubleDipper exhibits strong performance overall, we observe nuanced behavior on the SRT dataset, with performance gains varying across models (improvement for Gemini, Llama 3.1 8B, Mistral 7B and Mistral Nemo).
+This discrepancy is likely due to the fact that SRT demands a specific type of reasoning, where models must reason over both a rule (e.g., “If X is big and X is good then X is tall”) and dispersed facts (e.g., “Erin is Good” and “Erin is furry”) to determine whether a statement (e.g., “Erin is tall”) can be derived from the context.
+Finally, the Lost dataset highlights a specific characteristic of DoubleDipper: while the baseline’s known positional bias (Liu et al., 2024a) masks the average improvement for some models, DoubleDipper substantially boost performance when relevant information appear in the middle of the context (further elaborated in Result 3), demonstrating its efficacy in mitigating positional biases in long-context settings.
 Avg. 2Wiki MonoRel PIR SRT HotPotQA Lost MuSique
@@
 DoubleDipper 79.1 83.7 98.3 80.2 71.2 84.6 66.1 69.5
-Gemini Nano Zero-shot + Evidence Retrieval 68.1 76.3 77.6 71.6 66.0 74.1 55.8 55.2
-DoubleDipper 69.9 76.6 86.6 74.5 59.8 76.0 59.2 56.7
-Gemma 2B Zero-shot + Evidence Retrieval 39.1 57.6 49.4 44.3 18.9 53.1 14.0 36.4
-DoubleDipper 45.6 56.5 60.7 57.5 13.1 58.1 33.1 39.9
-Gemma 9B Zero-shot + Evidence Retrieval 61.9 76.7 69.3 60.3 43.2 76.5 51.5 55.9
+Gemma 2 9B Zero-shot + Evidence Retrieval 61.9 76.7 69.3 60.3 43.2 76.5 51.5 55.9
 DoubleDipper 57.0 74.2 52.5 59.0 23.1 78.4 55.3 56.8
+Gemma 2 27B Zero-shot + Evidence Retrieval 85.1 96.2 97.9 96.1 84.1 90.7 53.4 77.3
+DoubleDipper 83.8 97.5 85.4 97.6 74.9 93.0 59.7 78.8
 Llama 3.1 8B Zero-shot + Evidence Retrieval 61.7 53.2 86.5 66.9 67.8 63.2 41.1 53.5
 DoubleDipper 61.4 68.9 71.4 54.9 52.8 73.7 53.0 54.8
+Llama 3.1 70B Zero-shot + Evidence Retrieval 85.1 98.2 98 89.2 82.3 93.9 52.5 81.6
+DoubleDipper 87.8 98.6 100 92.8 83.8 96.5 61.6 81.6
 Mistral 7B (v0.3) Zero-shot + Evidence Retrieval 46.6 62.4 49.8 46.2 17.5 64.0 43.4 43.0
@@
 DoubleDipper 58.9 70.4 81.6 63.6 18.3 75.0 50.4 53.2
+Mistral-Nemo Zero-shot + Evidence Retrieval 69.7 91.1 85.0 76.9 35.9 88.8 39.8 70.3
+DoubleDipper 76.1 95.9 95.1 81.5 39.0 93.7 54.8 73.0
 Table 3: Performance (F1) of supporting paragraph(s) prediction.
-Figure 4: Performance (accuracy) of Gemma 2 9B, Mixtral 8x7B and Gemini Flash with and without DoubleDipper on our sample of the Lost-in-the-middle dataset (Liu et al., 2023) according to the position of the document that contains the answer.
-Result 2: Learning to retrieve the evidence(s) with DoubleDipper is more effective in commercial models than in open source.
+Result 2: Learning to retrieve the evidence(s) with DoubleDipper is more effective in commercial and large open source models.
 Table 3 presents the performance of the supporting paragraphs prediction for the Zero-shot + Evidence Retrieval and DoubleDipper on our evaluation set.
-For all commercial models and Gemma 2B, DoubleDipper predicts better the supporting paragraphs than in the zero-shot setting (+2.6 F1 for Gemini Pro, +3.4 F1 for Gemini Flash, +1.8 F1 for Gemini Nano and +6.5 F1 for Gemma 2B).
-On the other hand, DoubleDipper slightly hurts the performance of common open source models (e.g., -2.8 F1 for Mistral and -0.3 F1 for Llama 3.1).
-This difference is because DoubleDipper’s demonstrations are based on a single paragraph, where commercial models can generalize better and predict multiple evidences.
-Indeed, Gemini Pro predicts on average 2 evidences for the datasets 2Wiki, MuSique, PIR, MonoRel, SRT and HotPotQA whereas Gemma 9B predicts 1.2 evidences for each instance.
-In fact, the only dataset with a single evidence is “Lost” and almost all models highly benefit from DoubleDipper (e.g, +11.9 F1 for Llama 3.1, +12.2 F1 for Mistral 7B and +9.8 for Mixtral 7x8B).
+For all commercial models, Llama 3.1 70B and the recent Mistral-Nemo-Instruct-2407, DoubleDipper predicts better the supporting paragraphs than in the zero-shot setting (+2.6 F1 for Gemini Pro, +3.4 F1 for Gemini Flash, +2.7 F1 for Llama 3.1 70B and +6.4 F1 for Mistral-Nemo-Instruct-2407).
+Conversely, DoubleDipper slightly hurts the performance of common open source models (e.g., -2.8 F1 for Mistral, -0.3 F1 for Llama 3.1, -1.3 F1 for Gemma 27B, etc.).
+This discrepancy appears to stem from shortcut learning (Tang et al., 2023).
+Indeed, our demonstrations in DoubleDipper use a single evidence paragraph, and smaller models tend to overfit to this pattern, learning to retrieve only one passage even when multiple are needed.
+For example, under DoubleDipper, Gemma 2 9B retrieves only 1.2 paragraphs on average, compared to 2 in the “unconstrained” zero-shot setting.
+Larger models like Gemini Pro do not exhibit this behavior, correctly generalizing to predict an average of 2 evidence paragraphs even with single-paragraph demonstrations.
+This hypothesis is strongly supported by the “Lost” dataset, which requires only a single evidence paragraph.
+On this dataset, nearly all models, including smaller ones, benefit substantially from DoubleDipper (e.g., +12.2 F1 for Mistral 7B).
+Crucially, despite their suboptimal retrieval performance on multi-evidence tasks, these smaller models still produce better final answers with DoubleDipper than the zero-shot baseline, highlighting the net benefit of the approach.
 Result 3: DoubleDipper makes models more robust to the position of relevant information.
 Following Liu et al.
-(2023), Figure 4 shows the performance of Gemma 2 9B, Mixtral 8x7B and Gemini Flash for both the baseline and DoubleDipper on our sample of the Lost-of-the-middle dataset, according to the position of the document that contains the answer.
-See Appendix B for the performance curve of the other tested models, which show similar trends.
+(2023b), Figure 3 shows the performance of Gemma 2 9B, Mixtral 8x7B and Gemini Flash for both the baseline and DoubleDipper on our sample of the Lost-of-the-middle dataset, according to the position of the document that contains the answer.
+See Appendix D for the performance curve of the other tested models, which show similar trends.
 Overall, the performance curve for DoubleDipper consistently surpasses the baseline when the relevant information appears “in the middle” and sometimes also at the beginning and/or the end (e.g., Gemini Flash).
@@
 This reveals that beyond improving performance, DoubleDipper can make the model more robust to the position of the relevant document.
+Figure 3: Performance (accuracy) of Gemma 2 9B, Mixtral 8x7B and Gemini Flash with and without DoubleDipper on our sample of the Lost-in-the-middle dataset (Liu et al., 2023b) according to the position of the document that contains the answer.
 6 Ablation Studies
-Gemma 2B Self Gemini Pro
-Gemini Pro 70.4 71.6 71.6
-Gemini Nano 62.1 61.7 62.8
-Gemini Flash 66.1 67.5 68.1
-Gemma 2B 49.5 49.5 51.0
-Gemma 9B 61.2 62.0 63.5
-Llama 3.1 59.9 60.1 61.5
-Mistral v0.3 51.0 49.8 52.6
-Mixtral 52.2 49.8 54.4
-Table 5: Average performance of DoubleDipper with different models for generating the demonstrations. See Appendix C.2 for the results on each evaluation dataset.
-How many examples are needed?
-In Table 6, we explore the impact of varying $k$, the number of self-generated few-shot examples in DoubleDipper to 1, 3, 5, and 10.
-On average, a single demonstration already provides an improvement over the baseline.
-Adding 3 demonstrations adds another boost of 2 points, while increasing the number of demonstrations to 5 and 10 leads to a marginal improvement.
-This finding is in line with previous work (Brown et al., 2020a; Min et al., 2022b).
-We conclude that a small number of examples carries most of the benefit with our method, but given additional computation budget, adding more examples does carry additional minor benefit.
 $k=1$ $k=3$ $k=5$ $k=10$
@@
 Mixtral 7x8B (v0.1) 49.3 52.2 51.7 52.2
-Table 6: Average performance on our evaluation set with various numbers of self-generated few shot demonstrations ($k$) in DoubleDipper. See Appendix C.1 for the results on each evaluation dataset.
+Table 4: Average performance on our evaluation set with various numbers of self-generated few shot demonstrations ($k$) in DoubleDipper. See Appendix E.1 for the results on each evaluation dataset.
+How many examples are needed?
+In Table 4, we explore for some models the impact of varying $k$, the number of self-generated few-shot examples in DoubleDipper to 1, 3, 5, and 10.
+On average, a single demonstration already provides an improvement over the baseline.
+Three demonstrations adds another boost of 2 points, while increasing the number of demonstrations to 5 and 10 leads to a marginal improvement.
+This finding is in line with previous work (Brown et al., 2020a; Min et al., 2022).
+We conclude that a small number of examples carries most of the benefit with our method, but given additional computation budget, adding more examples does carry additional minor benefit.
+DoubleDipper without identification of supporting paragraphs
+To ablate the second core principle of DoubleDipper—the explicit identification of supporting paragraphs prior to answer generation—we prompt open-source models with self-generated few-shot examples consisting solely of question-answer pairs, without instructing the model to retrieve the relevant passage(s).
+These demonstrations may undermine DoubleDipper’s objective by encouraging models to produce answers without grounding them in the source text.
+The results confirms our hypothesis: removing evidence identification consistently degrades QA performance compared to the full DoubleDipper approach.
+When averaging results across models and datasets, this omission leads to a substantial performance drop from 54.8 to 46.6.
+Detailed results are provided in Appendix E.3, Table 10.
+Gemma 2B Self Gemini Pro
+Gemini Pro 70.4 71.6 71.6
+Gemini Flash 66.1 67.5 68.1
+Gemini Nano 62.1 61.7 62.8
+Gemma 2B 49.5 49.5 51.0
+Gemma 9B 61.2 62.0 63.5
+Llama 3.1 59.9 60.1 61.5
+Mistral v0.3 51.0 49.8 52.6
+Mixtral 52.2 49.8 54.4
+Table 5: Average performance of DoubleDipper with different models for generating the demonstrations. See Appendix E.2 for the results on each evaluation dataset.
 Investigating the effect of the few-shot generator
@@
 In the second experiment, we generate the demonstrations with the best LLM used in our experiments, namely Gemini Pro.
-The average results are reported in Table 5 and the performance for each evaluation dataset is presented in Appendix C.2.
+The average results are reported in Table 5 and the performance for each evaluation dataset is presented in Appendix E.2.
 The results show that generating the demonstrations with Gemma 2 or Self achieves similar performance, while Gemini Pro leads to a consistent increase in performance across models, indicating that future better models can improve further the performance.
-As mentioned in Section 3, DoubleDipper promotes also efficiency by adding to the original prompt only a few extra tokens, leading to a significantly cheaper inference than the traditional ICL.
-DoubleDipper without identification of supporting paragraphs
-To ablate the second principle in DoubleDipper, which is the explicit identification of the supporting paragraphs before generating the answer, we prompt the open source models with self-generated few shot examples that comprise only question-answer pairs (without instructing the model to retrieve the relevant passage(s)).
-For all models, the average QA performance of DoubleDipper without the evidence is lower than DoubleDipper with evidence identification.
-When averaging results across models and datasets, we report a substantial F1 drop from 54.8 to 46.6.
-Full results are available in Appendix C.3 in Table 9.
+Please refer to Appendix E.2 for additional ablations on the few-shot generation.
+Baseline ICL DoubleDipper
+Gemma 2 9B 44.0 51.0 61.2
+Gemma 2 27B 48.0 54.9 64.2
+Llama 3.1 8B 37.2 40.7 59.9
+Llama 3.1 70B 67.5 65.4 72.9
+Mistral v0.3 37.4 42.0 51.0
+Mixtral v0.1 42.6 45.1 52.2
+GPT 4o mini 51.3 56.1 60.8
+Table 6: Comparison of traditional In-Context Learning (ICL) where each demonstration example comprises a full text, a question and an answer from an external dataset to DoubleDipper where the demonstrations contain only question-answer pairs, automatically generated on the same input text.
+DoubleDipper vs. Traditional ICL
+Another alternative to use ICL in practice is to preprend each QA prompt by a fixed set of QA demonstrations, each composed of a context, a question and an answer.
+Although the demonstrations are not necessarily from the same distribution as each user query, this common practice is helpful for task recognition and for an overview of the overall format (Min et al., 2022; Pan et al., 2023).
+For the demonstrations, we randomly selected 3 examples from the SQuAD 2.0 dataset (Rajpurkar et al., 2018).
+We compare the average results of the baseline, ICL and DoubleDipper in Table 6 and report full results in Appendix E.4 in Table 11.
+While ICL is effective and outperforms the baseline, DoubleDipper provides an additional performance boost of 9.5 points on average.
 Qualitative analysis: Correctness of the generated QA pairs
@@
 7 Conclusion
-We develop DoubleDipper, a straightforward method for enhancing the performance of Question Answering with long context and providing attribution to the relevant paragraph(s) in the input.
-By recycling the input context to generate the few shot examples, each demonstration includes solely a question, an answer and a pointer to the relevant paragraph, without a separate context, thus effectively addressing the challenging of In-Context-Learning with long context.
-Experimental results show that our approach substantially ourperforms the baselines in various QA settings, including distractor passages in the input, True/False questions and multi-hop QA.
+We introduce DoubleDipper, a simple method for enhancing the performance of LLMs with long context.
+By recycling the input context to generate the demonstrations, DoubleDipper successfully addresses the practical challenges of ICL with long-context and outperforms multiple baselines in various QA settings, including distractor passages in the input, True/False questions and multi-hop QA.
 8 Limitations
-One notable limitation of our approach is the extended inference time required for generating question-answer pairs.
-Future research could mitigate this issue by developing smaller, specialized models specifically tailored for QA generation.
-Additionally, our evaluation set is constrained to instances that are solely in English and range between 1,000 to 4,000 tokens.
-Expanding the diversity of languages and token ranges could enhance the robustness and applicability of our findings.
-Lastly, although we employ a strategy of randomly sampling $k$ paragraphs from the input to ensure the model engages with varied segments of the text, we did not optimize the selection of these paragraphs.
-Future work could explore more strategic methods for paragraph selection to potentially enhance the efficacy and relevance of the generated examples.
-Appendix A Prompts
+Our work has several limitations.
+First, our current work focuses on multiple variants of QA tasks (distractor passages, True/False and multi-hop), where the demonstrations teach the models to identify specific evidence paragraph(s) and extract answers.
+Future research can extend our work to other QA settings (e.g., information seeking) and additional tasks (e.g., summarization).
+Second, our evaluation set is constrained to instances that are solely in English and range between 1,000 to 4,000 tokens.
+While this demonstrates the method’s effectiveness, its scalability and performance on much longer contexts (e.g., 100k+ tokens) and in multilingual settings remain open questions.
+Finally, while DoubleDipper is significantly more token-efficient than traditional ICL, the initial step of generating demonstrations introduces a computational and latency overhead compared to a zero-shot baseline.
+This presents a trade-off between inference cost and the substantial performance gains our method provides.
+Appendix A Appendix
+Appendix B Prompts
 Figure 4 shows the zero-shot prompt we use for generating the question-answer pairs in DoubleDipper.
@@
 Figure 4: Template prompt for generating the QA pairs in DoubleDipper.
-Appendix B Lost-in-the-middle
+Appendix C Additional Results
+Table 7 shows the QA performance of the baseline, Zero-shot + Evidence Retrieval and DoubleDipper on our evaluation set when applied to Gemma 2B, Gemini Nano and GPT-4o-mini. Her
+Avg. 2Wiki MonoRel PIR SRT HotPotQA Lost MuSique
+Gemini Nano (vanilla) 41.6 10.8 72.2 66.8 55.4 21.3 59.6 5.2
+   Zero-shot + Evidence Retrieval 56.5 32.0 82.4 82.4 56.4 56.7 60.5 25.2
+   DoubleDipper 62.1 40.6 86.6 95.4 56.2 65.1 60.5 30.4
+GPT-4o-mini (vanilla) 51.3 19.0 78.6 89.6 71.6 23.3 67.8 9.3
+   Zero-shot + Evidence Retrieval 53.8 19.4 89.6 93.0 63.2 26.2 67.9 17.6
+   DoubleDipper 60.8 29.8 94.8 96.4 61.6 53.7 64.7 24.5
+Gemma 2 2B (v2) (vanilla) 38.6 8.9 71.8 68.6 51.2 13.3 49.6 6.5
+   Zero-shot + Evidence Retrieval 42.0 22.3 66.8 70.6 40.2 30.6 47.6 16.2
+   DoubleDipper 49.5 23.7 85.8 81.6 50.0 39.9 46.7 18.8
+Mixtral 7x8B (v0.1) (vanilla) 42.6 13.7 73.0 66.2 51.0 18.2 67.7 8.4
+   Zero-shot + Evidence Retrieval 47.4 18.8 81.8 73.6 50.6 26.3 67.9 13.1
+   DoubleDipper 52.2 22.3 91.8 86.0 47.8 35.1 66.6 16.0
+Table 7: Accuracy of the QA task for the vanilla baseline (prompting the model to only answer the question), Zero-shot + Evidence Retrieval (prompting the model to explicitly identify the relevant passage(s) before generating the answer) and DoubleDipper with 3 demonstrations generated by Gemma 2 2B.
+Appendix D Lost-in-the-middle
 Figure 5 shows the QA accuracy of the models Gemma 2 2B, Mistral 7B, Gemini Nano and Gemini Pro on our subset of the “Lost-in-the-middle” dataset.
-Figure 6: Performance (accuracy) of Gemma 2 2B, Mistral 7B, Gemini Nano and Gemini Pro with and without DoubleDipper on our sample of the Lost-in-the-middle dataset (Liu et al., 2023) according to the position of the document that contains the answer.
-Appendix C Analysis
-C.1 Impact of the Number of Demonstrations in DoubleDipper
+Figure 5: Performance (accuracy) of Gemma 2 2B, Mistral 7B, Gemini Nano and Gemini Pro with and without DoubleDipper on our sample of the Lost-in-the-middle dataset (Liu et al., 2023b) according to the position of the document that contains the answer.
+Appendix E Analysis
+E.1 Impact of the Number of Demonstrations in DoubleDipper
 Avg. 2Wiki MonoRel PIR SRT HotPotQA Lost MuSique
@@
 Figure 6: Performance (accuracy) of several models with DoubleDipper according to the number of self-generated demonstrations in the prompt.
-C.2 Impact of the few-shot generator
+E.2 Impact of the few-shot generator
 Table 9 presents the detailed QA performance of all models with different models for generating DoubleDipper’s demonstrations.
@@
 Table 9: Performance of DoubleDipper according to the model for generating the demonstrations (Gemma 2B, Self or Gemini Pro).
-C.3 Impact of the identification of supporting paragraphs in the QA generation
+Can we generate demonstrations with smaller LLMs?
+We next study whether smaller, more efficient models could be used to generate demonstrations for DoubleDipper.
+To do this, we created demonstrations using two smaller LLMs, Gemma 3 1B (Kamath et al., 2025) and Qwen 2.5 0.5B (Bai et al., 2023a), and provid them to a subset of our evaluation models (Llama 3.1, Gemma 2 9B, Gemma 2 27B, and Mistral Nemo).
+The results show that DoubleDipper maintains its advantage over baselines, exhibiting only a modest performance drop of less than 2 points compared to using demonstrations from the larger Gemma 2 2B.
+This finding suggests that DoubleDipper remains effective even when its demonstration generation component is replaced with more lightweight models.
+Can DoubleDipper benefit from incorrect demonstrations?
+To answer this question, we prompt Gemma 2 2B to generate a question and an incorrect answer and provide these demonstrations to a sample of our tested models (Llama 3.1, Gemma 2 9B, Gemma 2 27B and Mistral Nemo).
+The results are mixed and not conclusive: some models barely benefit or suffer from incorrect demonstrations (+1.6 for Gemma 2 9B, -0.4 for Gemma 2 27B), while others somehow benefit from incorrect demonstrations (+4.8 for Llama 3.1 8B and +5.1 for Mistral Nemo over the baseline).
+Critically, however, all models still perform substantially worse than DoubleDipper with correct demonstrations.
+A possible explanation for the unexpected gains is that even incorrect examples provide useful structural guidance for the task format, a phenomenon observed in prior work (Min et al., 2022).
+E.3 Impact of the identification of supporting paragraphs in the QA generation
 Table 10 compares the performance of DoubleDipper to DoubleDipper without evidence identification.
@@
 Table 10: Performance of DoubleDipper and DoubleDipper without instructing the models to retrieve the evidence (QA only) on the QA datasets.
+E.4 In-Context-Learning
+Table 11 presents the results of our tested models when prepended with three in-context demonstrations, taken from the Squad 2.0 dataset (Rajpurkar et al., 2018).
+Avg. 2Wiki MonoRel PIR SRT HotPotQA Lost MuSique
+Gemma 2 2B 40.4 11.5 69.4 59.8 50.6 33.1 48.8 9.7
+Gemma 2 9B 51.0 21.7 76.8 78.8 53.6 49.1 60.0 17.3
+Gemma 2 27B 54.9 21.3 85.0 76.6 58.2 53.3 61.5 28.3
+Llama 3.1 8B 40.7 16.4 59.6 53.0 45.6 36.8 62.5 11.0
+Llama 3.1 70B 65.4 38.1 93.0 91.4 75.4 59.7 67.6 32.4
+Mistral v0.3 42.0 25.5 56.4 52.8 50.2 42.7 58.2 8.3
+Mixtral v0.1 45.1 19.6 73.6 65.4 50.4 30.7 66.6 9.6
+GPT 4o mini 56.1 29.7 77.4 86.8 65.2 49.4 65.6 18.5
+Table 11: Accuracy of the QA task for the ICL experiment with 3 fixed demonstrations prepended to each instance.